### v1

# Conformalized Late Fusion Multi-View Learning

Abstract
Uncertainty quantification for multi-view learning is motivated by the increasing use of multi-view data in scientific problems. A common variant of multi-view learning is late fusion: train separate predictors on individual views and combine them after single-view predictions are available. Existing methods for uncertainty quantification for late fusion often rely on undesirable distributional assumptions for validity. Conformal prediction is one approach that avoids such distributional assumptions. However, naively applying conformal prediction to late-stage fusion pipelines often produces overly conservative and uninformative prediction regions, limiting its downstream utility. We propose a novel methodology, Multi-View Conformal Prediction (MVCP), where conformal prediction is instead performed separately on the single-view predictors and only fused subsequently. Our framework extends the standard scalar formulation of a score function to a multivariate score that produces more efficient downstream prediction regions in both classification and regression settings. We then demonstrate that such improvements can be realized in methods built atop conformalized regressors, specifically in robust predict-then-optimize pipelines.
1 Introduction
Multi-view machine learning is becoming increasingly common with the advent of multi-modal language models and conditional generative models [1, 2, 3, 4, 5]. In addition, such techniques are being increasingly used in more safety-critical settings, such as medicine [6, 7, 8, 9] and robotics [10, 11, 12, 13]. These use cases motivate the development of uncertainty quantification methods for multi-view learning.
The most common division of multi-view learning is along the axis of when fusion of the multiple views is performed. We can either learn a fused data representation or we can fuse the predictions from the separate views. The latter approach is referred to as “late fusion”. It is desirable in limited resource settings, where the separately trained models can be combined in a light-weight fashion for downstream tasks [14, 15, 16, 17]. Increasing interest is now being placed on quantifying uncertainty for such models [18, 19, 20, 21, 22]. Such methods, however, rely on distributional assumptions to provide coverage guarantees or are computationally intractable, limiting their practical utility. These limitations, therefore, demand a more universal, yet informative, uncertainty quantification method.
One such method for performing distribution-free uncertainty quantification is conformal prediction, which provides a principled framework for producing distribution-free prediction regions with marginal frequentist coverage guarantees [23, 24]. By using conformal prediction on a user-defined score function $s(x,y)$ and obtaining an empirical $1-\alpha$ quantile $\widehat{q}(\alpha)$ of $s(x,y)$ over a calibration set $\mathcal{D}_{C}$, prediction regions $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}(\alpha)\}$ attain marginal coverage guarantees. While calibration is guaranteed from this procedure, predictive efficiency, quantified as the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
We extend the conformal prediction line of work to enable its practical use in late-stage multi-view prediction problems. In particular, we propose a multivariate extension to the score function notion introduced in traditional conformal prediction, with a corresponding extension of quantiles to “quantile envelopes.” We demonstrate that this formulation retains the desired distribution-free coverage guarantees typical of standard conformal prediction and that the resulting prediction regions can be used efficiently in both classification and regression settings. Our contributions are:
Figure 1: MVCP provides a principled extension to the standard conformal prediction pipeline by leveraging ideas from higher-dimensional quantile regression to define quantile envelopes $\widehat{\mathcal{Q}}$ instead of purely scalar quantiles $\widehat{q}$. It does so by evaluating a collection of score functions (here $s_{1}$ and $s_{2}$) over the calibration dataset to define $\mathcal{S}$, finding quantiles $\{\widehat{q}_{m}\}$ over a set of projection directions $\{u_{m}\}$, and taking $\widehat{\mathcal{Q}}$ to be the intersection of the resulting half-planes $H(u_{m},\widehat{q}_{m})$. These quantile envelopes result in more informative prediction regions that can be used in downstream tasks.
  • Providing a multivariate extension to conformal prediction score functions (MVCP) that retains coverage guarantees and can be applied to late-stage multi-view prediction problems.
  • Demonstrating how the resulting prediction regions can be efficiently leveraged for downstream classification and regression tasks.
  • Demonstrating the empirical performance of the MVCP framework across a suite of tasks.
2 Background
2.1 Conformal Prediction
Coverage guarantees of uncertainty quantification methods generally rely on distributional assumptions, often via asymptotics or explicit model specification. To alleviate the need for such restrictive assumptions, interest in finite-sample, distribution-free uncertainty quantification methods has risen. Conformal prediction is one such method [23, 24]. Due to its lack of structural assumptions, conformal prediction pairs well with the increasingly common black-box style of modeling.
Conformal prediction serves as a wrapper around such predictors, producing prediction regions $\mathcal{C}(x)$ that have formal guarantees of the form $\mathcal{P}_{X,Y}(Y\notin\mathcal{C}(X))\leq\alpha$ for some prespecified level $\alpha$. To achieve this, “split conformal” partitions the dataset $\mathcal{D}$ into a training set $\mathcal{D}_{T}$ and a calibration set $\mathcal{D}_{C}$. The former plays the role traditionally played by $\mathcal{D}$, namely serving as data used to fit $\widehat{f}$. The latter is used much in the way a validation set is typically employed in practice. Formally, users of conformal prediction must design a “score function” $s(x,y)$, which should quantify “test error,” often in a domain-specific manner. For instance, a simple score function for a regression setting would be $s(x,y)=\|\hat{f}(x)-y\|$. This score function is then evaluated across the calibration set to define $\mathcal{S}=\{s(x,y)\mid(x,y)\in\mathcal{D}_{C}\}$. For a desired coverage of $1-\alpha$, we then take $\widehat{q}$ to be the $\lceil(|\mathcal{D}_{C}|+1)(1-\alpha)\rceil/|\mathcal{D}_{C}|$ quantile of $\mathcal{S}$, with which prediction regions for future test queries $x$ can be defined as $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}\}$. Under the exchangeability of the score of a test point $s(x^{\prime},y^{\prime})$ with $\mathcal{S}$, we have the desired finite-sample probabilistic guarantee that $1-\alpha\leq\mathcal{P}_{X,Y}(Y\in\mathcal{C}(X))$.
While this guarantee holds for any $s(x,y)$, the informativeness of the resulting prediction regions, formally defined as the inverse expected Lebesgue measure across $X$, i.e. $\left(\mathbb{E}[\mathcal{L}(\mathcal{C}(X))]\right)^{-1}$, is intimately tied to its specification [24]. Thus, much of the practical challenge of conformal prediction relates to choosing a score function that retains coverage while minimizing prediction region size.
2.2 Predict-Then-Optimize
Prediction regions produced from conformalized predictors are often leveraged differently across classification and regression settings. In the case of classification, “regions” simply constitute a subset of the label space, making its direct use by end users straightforward, as demonstrated in works such as [25]. In high-dimensional regression settings, however, prediction regions become harder to use directly; for this reason, recent works have started shifting focus to using them in their implicit forms.
One such application is [26], where conformal prediction was leveraged in a predict-then-optimize setting. As the name suggests, predict-then-optimize problems are two-stage problems, which take observed contextual information $x$ and predict the parametric specification of a downstream problem of interest $\widehat{c}:=g(x)$ with some trained predictor $g$. The final result is then a decision made with this specification, $w^{*}:=\min_{w}f(w,\widehat{c})$. An example of such a setting is if an optimal labor allocation $w^{*}$ is sought based on predicted demand $\widehat{c}$ from observed transactions $x$ in a food delivery platform.
While the predicted $\widehat{c}$ is often simply trusted in the downstream optimization, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to highly suboptimal decision-making. For this reason, recent interest has been placed on studying a “robust” alternative to this formulation [27, 28, 29]. Following this line of work, [26] proposed studying the following formulation: $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with the uncertainty region $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$. The primary consequence of doing so was having a probabilistic guarantee on the suboptimality gap. That is, denoting the suboptimality gap by $\Delta(x,c):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})-\min_{w}f(w,c)$, where $c$ is the true parameter corresponding to $x$, they established that leveraging conformal uncertainty regions guarantees $\mathcal{P}_{X,C}\left(0\leq\Delta(X,C)\leq L\mathrm{\ diam}(\mathcal{C}(X))\right)\geq 1-\alpha$ if $f(w,c)$ is an objective function that is convex-concave and $L$-Lipschitz in $c$ for any fixed $w$.
2.3 Multi-View Learning
We provide a brief overview of multi-view learning herein, deferring to recent surveys of the field for a more comprehensive review thereof [30]. In general, the focus of multi-view machine learning is to solve prediction problems when different, often non-overlapping, representations of the same underlying “object” are observed. Methods in this space bifurcate between those defined together with some downstream supervised task and those seeking to find an abstract “fused” representation that lends itself naturally for arbitrary downstream use cases. We focus herein on the former category. For instance, one may wish to classify a captioned image, where the image and caption represent two distinct “views” of the same underlying abstract object in this case.
Formally, therefore, we consider the multi-view prediction problem as seeking a map $\mathcal{X}\rightarrow\mathcal{Y}$ with $\mathcal{X}=\mathcal{X}_{1}\cup...\cup\mathcal{X}_{K}$. As previously mentioned, fusion approaches are generally either early-stage or late-stage. In the former, we seek to learn some common, often abstract, space into which these separate views can be mapped for downstream prediction, namely seeking a map $g_{\phi}$ that maps $g_{\phi}:\mathcal{X}_{1}\times...\times\mathcal{X}_{K}\rightarrow\mathcal{Z}$. Such a fusion map is then assumed to be usable for the downstream prediction task, namely through composition with $f_{\theta}:\mathcal{Z}\rightarrow\mathcal{Y}$ such that the overall prediction model can be defined as $f\circ g$. Traditionally, $g$ was hand-crafted, leveraging domain-specific insights. However, increasingly, $f$ and $g$ are simply subsets of a neural network trained end-to-end via gradient descent, such as in [31, 32, 33, 34, 35, 36].
As discussed previously, this fusion approach is not universally desirable in all circumstances though, such as when models for individual views have already been curated after computationally intensive training and hyperparameter tuning [14, 15, 16, 17]. In such cases, late-stage fusion approaches can be preferred, where separate predictors $f_{i}:\mathcal{X}_{i}\rightarrow\mathcal{Y}$ have been trained and are fused via a post-hoc algorithm $\mathcal{F}:\mathcal{Y}\times...\times\mathcal{Y}\rightarrow\mathcal{Y}$. Approaches to this include decision averaging [37, 38], decision weighting [18, 39, 40], and decision voting [41]. We focus herein on such late-stage fusion.
2.4 Related Works
Work in the conformal prediction literature has yet to leverage the generalized notions of quantiles considered herein. Generalizations of quantiles, while introduced first in the context of conformal prediction herein, have a long history in traditional statistics, with recent surveys on approaches available in [42, 43]. We briefly summarize the relevant pieces of this literature below.
Unlike univariate data, multivariate data does not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering of points in higher dimensional spaces. To account for this, one class of multivariate quantiles, known as “direction quantiles,” prescribes directions along which data points can be compared [44, 45, 46]. That is, for some direction $u\in\mathcal{S}^{n-1}$, points $x\in\mathbb{R}^{n}$ are projected, after which the standard notion of a quantile can be leveraged.
The notion of the $\alpha$ quantile for a random variable $X\in\mathcal{X}$ clearly then depends on this choice of direction $u$, which we denote by $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$. When there is no ambiguity, we just denote it as $Q(\alpha,u)$. In cases where domain knowledge assigns an unambiguous notion of multivariate ordering, such knowledge can be encoded into the choice of $u$. In other cases, however, a more agnostic approach is typically employed: rather than choosing any particular direction, all directions are considered simultaneously. Formally, for any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$. To extend this to consider all directions simultaneously, we then define the quantile envelope to be
$D(\alpha)=\bigcap_{u\in\mathcal{S}^{n-1}}H(u,Q(\alpha,u)).$ (1)
To the best of our knowledge, the only work leveraging conformal prediction in the context of multivariate quantiles is [47]. This work, however, focuses on a fundamentally different problem than that considered herein. In particular, they focused on leveraging conformal prediction as a correction step to the output of directional quantile regression with the intended use case being in constructing calibrated prediction regions for multivariate regression problems. This work, therefore, was a natural extension to the now classical application of conformal prediction for the correction of univariate quantile regression. Our proposed method, however, does the converse: it leverages directional quantiles within the score space to extend conformal prediction itself.
3 Method
We now introduce a procedure for uncertainty quantification for late-stage fusion. We discuss the naive approach in Section 3.1 and its extension along with the associated coverage guarantees in Section 3.2. We then discuss the downstream use of the resulting prediction regions in Section 3.3.
3.1 Naive Approach
We now consider the setting described in the background section, where $K$ predictors $f_{1},...,f_{K}$ are trained on the separate views with a late fusion algorithm $\mathcal{F}$, which we assume to be fixed throughout this presentation. Note that the validity of the results presented below is independent of the predictive quality of such $f_{i}$. The efficiency, however, is highly dependent on it. In the standard application of conformal prediction, a score function $s(\mathcal{F}(x),y)$ is defined. For instance, if we assume $\mathcal{Y}$ is some metric space with metric $d$, as in a regression task, a naive choice of score function would be
$s(x,y)=d(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y).$ (2)
Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ is then marginally calibrated. Ideally, however, prediction regions $\mathcal{C}(x)$ should have the quality that, if a particular predictor has less uncertainty in its predictions, as is frequently true of settings leveraging late-stage fusion, upon routing to that predictor, the corresponding size of the prediction region should be smaller than if it had been routed to a different predictor. While the framing of Equation 2 does, in principle, support this property, it ultimately relies on defining an uncertainty-aware late fusion algorithm $\mathcal{F}$. In its typical form, however, $\mathcal{F}$ simply takes point predictions $f_{1}(x),...,f_{K}(x)$ in as input, meaning any uncertainty-awareness would need to be baked in a priori through domain knowledge of the uncertainties of the predictors $f_{1},...,f_{K}$, which can seldom be specified precisely, in turn sacrificing the downstream predictive efficiency of $\mathcal{C}(x)$.
For this reason, we seek to produce a data-driven approach of obtaining the uncertainties of the predictors by instead fusing the conformalized predictors. In this case, we first define a score function $s_{i}$ for each predictor $f_{i}$. For any $x$, therefore, we now have a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$. A naive approach would then involve leveraging standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$. As in the case of conformalizing the fused predictor directly, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties. We instead wish to provide a data-adaptive pipeline to automatically produce such a $g$. Formally, denoting by $\mathcal{C}(g):\mathcal{X}\rightarrow\mathcal{Y}$ the conformalized predictor resulting from calibrating with the score fusion function $g$ in some function class $\mathcal{G}$, the goal can be formulated as seeking $g^{*}:=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{X}\left[\mathcal{L}\left(\mathcal{C}\left(g\right)\left(X\right)\right)\right]$. There is, however, no approach to solving this abstract formulation, since even the mere evaluation of $\mathbb{E}_{X}[\mathcal{L}(\mathcal{C}(g)(X))]$ for a fixed $g$ is intractable.
3.2 MVCP: Score Quantile Envelope
Since finding $g^{*}$ exactly is intractable, we propose a data-adaptive approach that leverages the notion of quantile envelopes from Section 2.4 over the scores $s\in\mathbb{R}^{K}$. Importantly, for the remainder of this discussion, we assume the score functions are non-negative, i.e., $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which is typically the case as the score serves as a generalization of the classical idea of a residual. Intuitively, the coverage guarantee in this multivariate extension requires us to define a pre-ordering $\lesssim$ over $\mathbb{R}^{K}$, where we only lose the antisymmetric axiom of a total ordering. Upon defining such a pre-ordering, the coverage guarantee follows in using the appropriate quantile under this modified sorting.
Crucially, however, the problem of choosing this pre-ordering closely parallels that of choosing $g^{*}$, where a poorly chosen pre-ordering will result in overly conservative quantile envelopes and, hence, prediction regions. For instance, using a lexicographical ordering $\lesssim_{\mathrm{Lex}}$ will result in axis-aligned hyper-rectangular quantile regions and similarly a norm-based ordering $\lesssim_{\mathrm{Norm}}$, defined as $s_{1}\lesssim_{\mathrm{Norm}}s_{2}\iff||s_{1}||_{2}\leq||s_{2}||_{2}$, in hyperspherical regions. Rather than explicitly prescribing a pre-ordering, $\lesssim$ can be implicitly defined by prescribing an indexed family of nested sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, such that $\mathcal{A}_{t_{1}}\subset\mathcal{A}_{t_{2}}$ for $t_{1}\leq t_{2}$. We then define $\lesssim$ by stating $s_{1}\lesssim s_{2}$ if $\forall t$, $s_{2}\in\mathcal{A}_{t}\implies s_{1}\in\mathcal{A}_{t}$.
While not strictly necessary, we restrict ourselves to the case where $\mathcal{A}_{t}$ is a family of sets defined by a fixed “base contour” that scales outward with $t$. Under this restriction, performing a comparison of $s_{1},s_{2}\in\mathbb{R}^{K}$ in the aforementioned fashion, i.e. checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the $t$ for which $\mathcal{A}_{t}$ “intersects” $s$. Notably, $t(s)$ is precisely the aforementioned data-driven score fusion function $g(s)$ of interest. Defining a data-adaptive $g(s)$, therefore, reduces to having a data-driven approach for defining the base contour of $\mathcal{A}_{t}$. We additionally restrict this base contour to be a convex set; if such a set were permitted to be nonconvex, computing $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$ would potentially be computationally expensive, increasing the difficulty in defining the resulting prediction regions. As demonstrated in Section 3.3, many downstream applications require such efficient computation and, hence, benefit from this restriction to convex sets.
To have tight quantile regions, we formally wish for the pre-ordering to have the property that the quantile region given by $\widetilde{\mathcal{Q}}:=\{s\mid s\lesssim\widehat{q}\}$, where $\widehat{q}$ is the $1-\alpha$ empirical quantile of $\mathcal{S}_{C}$ under $\lesssim$, should have minimal Lebesgue measure. Notably, by construction, such a $\widetilde{\mathcal{Q}}$ will capture $1-\alpha$ points of $\mathcal{S}_{C}$. The problem of discovering an optimal pre-ordering can, thus, be equivalently stated as seeking to define the base contour of $\mathcal{A}_{t}$ to match that of the tightest $1-\alpha$ convex cover of $\mathcal{S}_{C}$.
This final reframing, therefore, naturally motivates selecting the base contour as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$. Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangebility of its points with future test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$. The goal, then, follows as seeking to define the base contour as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$. For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute the final $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$. Such a split is predicated on the assumption that the $1-\alpha$ quantile envelope defined over $\mathcal{S}_{C}^{(1)}$ resembles that of $\mathcal{S}_{C}^{(2)}$, implying the $|\mathcal{S}_{C}^{(1)}|$ should be sufficiently large as to capture this structure accurately.
We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$, starting by selecting the projection directions $\{u_{m}\}$. The restriction of $s$ to the positive orthant induces a natural modification of Equation 1, namely where $u\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$ instead of $\mathcal{S}^{K-1}$. Exact solution of finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [48]. If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values of $v$.
We now wish to define the quantile thresholds $\{\widetilde{q}_{m}\}$ for the selected directions to optimally capture $1-\alpha$ of $\mathcal{S}_{C}^{(1)}$. Naively taking the $1-\alpha$ quantile per projection direction $u_{m}$ results in joint coverage by $\widetilde{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$ of $\mathcal{S}^{(1)}_{\mathcal{C}}$ to be $<1-\alpha$. To combat this, a straightforward fix is to replace the $1-\alpha$ quantile per direction instead with its Bonferroni-corrected $1-\alpha/M$ quantile. While valid, this approach produces overly conservative prediction regions. We, therefore, instead tune a separate $\beta\in(\alpha/M,\alpha)$ parameter via binary search, finding the maximum $\beta^{*}$ such that using the $\beta^{*}$ quantile per direction provides the overall desired coverage, i.e. $|\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m}(1-\beta^{*}))\cap\mathcal{S}^{(1)}_{\mathcal{C}}|/N_{\mathcal{C}_{1}}\in(1-\alpha,1-\alpha+\epsilon)$ for some fixed, small $\epsilon>0$. With this choice of $\{(u_{m},\widetilde{q}_{m})\}$, we have a defined pre-ordering, whose coverage guarantees are formally stated below and proven in Appendix B.
Theorem 3.1.
Suppose $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$ are exchangeable. Assume further that $K$ maps $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined. Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\}_{(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}}$. For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ induced by a collection of nested sets $\{\mathcal{A}_{t}\}_{t\geq 0}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$. Then, denoting $\mathcal{C}(X):=\{y:s(X,y)\in\mathcal{Q}(\alpha)\}$, $\mathcal{P}_{X,Y}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha$.
To compute $\widehat{q}$, for each $s\in\mathcal{S}^{(2)}_{\mathcal{C}}$, we find $t^{*}(s)$, defined to be $\min\{t\in\mathbb{R}:s\in\bigcap_{m=1}^{M}H(u_{m},t\widetilde{q}_{m})\}$. This can be efficiently computed as $t^{*}(s)=\max_{m=1,...,M}(u_{m}^{\top}s/\widetilde{q}_{m})$. Denoting the $\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil$-th largest $t^{*}(s)$ as $\widehat{t}$, $\widehat{q}_{m}:=\widehat{t}\widetilde{q}_{m}$ and $\widehat{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widehat{q}_{m})$. Note that, if the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$.
Importantly, while the aforementioned procedure will necessarily result in convex regions $\widehat{\mathcal{Q}}(\alpha)$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, discussed more fully in Section 3.3. Such flexibility is necessary, as many studies have demonstrated the need for nonconvex prediction regions for downstream utility, such as [26, 49, 47]. However, it is unsurprising such flexibility exists, as even a single scalar score $s_{1}(x,y)$ can produce nonconvex prediction regions.
Another point to emphasis is that this procedure does not require a fusion algorithm to be defined: this is a key conceptual separation between MVCP and traditional conformal prediction. In the classic approach, a fused predictor is conformalized; here, instead, the conformalized predictors are fused. As a result, there is no notion of a point prediction for MVCP: it only returns prediction regions. This is fundamentally unlike the traditional case, where a point prediction $\widehat{y}:=\mathcal{F}(x)$ is “wrapped around” by a surrounding prediction region. We present the full algorithm in Algorithm 1 and an accompanying visual walkthrough of the procedure in Appendix A.
Algorithm 1 MVCP: $\textsc{UnifHypersphere}(K)$ is a subroutine that samples $\sim\mathrm{Unif}(\mathcal{S}^{K-1})$.
1:Inputs: Score functions $s_{1},...,s_{K}:\mathcal{X}\rightarrow\mathcal{Y}$, Calibration set $\mathcal{D_{C}}$, Desired coverage $1-\alpha$
$[\beta_{\mathrm{lo}},\beta_{\mathrm{hi}}]\leftarrow[\alpha/M,\alpha],\widehat{\mathcal{Q}}\leftarrow\emptyset,\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{1}(x_{i},y_{i}),...,s_{K}(x_{i},y_{i}))\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}}$
3:$\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
4:while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
5:     $\beta\leftarrow\frac{\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}}}{2}$
6:     $\left\{\widetilde{q}_{m}\leftarrow\frac{\lceil(N_{\mathcal{C}_{1}}+1)(1-\beta)\rceil}{N_{\mathcal{C}_{1}}}\text{ quantile of }\{u_{m}^{\top}s_{i}\}_{i=1}^{N_{\mathcal{C}_{1}}}\right\}_{m=1}^{M}$
7:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
8:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$
9:     else $\beta_{\mathrm{hi}}\leftarrow\beta$
10:$\widehat{t}\leftarrow\frac{\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil}{N_{\mathcal{C}_{2}}}\text{ quantile of }\{\max_{m=1,...,M}(u_{m}^{\top}s_{i}/\widetilde{q}_{m})\}_{i=1}^{N_{\mathcal{C}_{2}}}$
11:Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
3.3 MVCP: Prediction Region Definition
With this generalization of the score function, a natural question is how to leverage the resulting prediction regions $\mathcal{C}(x)$. For both classification and regression, $\mathcal{C}(x)=\{y\mid s(x,y)\in\widehat{\mathcal{Q}}\}$, or $\mathcal{C}(x)=\bigcap_{m=1}^{M}\mathcal{C}_{m}(x)$ where $\mathcal{C}_{m}(x):=\{y\mid u_{m}^{\top}s(x,y)\leq\widehat{q}_{m}\}$.
In the classification setting, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward. For a fixed $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the defining hyperplanes after projection.
In the case of regression, however, the prediction region cannot be explicitly constructed in the general case, since $\mathcal{Y}$ contains uncountably many elements. In fact, explicit construction is generally not of interest for downstream regression applications. Instead, prediction regions in regression settings are often simply used in their implicit form. We, therefore, focus on one particular application, namely that of [26] discussed in Section 2.2, and demonstrate the MVCP prediction regions can be leveraged in their framework. Such an extension has natural applications to the settings discussed in this original study. For instance, the authors demonstrated their method in a robust traffic routing setting with $c$ being predicted traffic from a probabilistic weather model $q(C\mid X)$ for weather covariates $X$. A multiview application, therefore, naturally emerges with having separately tailored predictive models, such as another $q_{2}(C\mid X_{2})$ predicting traffic based on historical traffic trends.
The original study demonstrated that solving this problem in a computationally efficient manner is feasible by leveraging Danskin’s Theorem so long as $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ is efficiently computable for any fixed $w$. We focus on demonstrating that this remains the case for MVCP, specifically considering the case where individual view score functions take the form of a generalization of the “GPCP” score from [26]. Let $q_{k}(C\mid X_{k})$ represent a generative predictor for view $k$ and $\{\widehat{c}_{jk}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X_{k})$ be drawn samples. Note that $J_{k}$ need not be constant across $k$. Then, we define the score to be
$s_{k}(x,c)=\min_{j}\left[g_{k}\left(\widehat{c}_{jk},c\right)\right],$ (3)
where $g_{k}(\widehat{c},c)$ is any function that is convex in $c$ for a fixed $\widehat{c}$. Notably this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X_{k})=\delta(f_{k}(X_{k}))$. To now compute $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, we first let $\vec{j}\in\mathcal{J}=\{j_{1},...,j_{K}\}$ be an indexing tuple, where each $j_{k}\in\{1,...,J_{k}\}$. That is, each $\vec{j}$ is a vector that “selects” one sample per view. Notably then, the projection $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)$ is convex in $c$, since the projection directions are all restricted to $\mathcal{S}_{+}^{K-1}$. As a result, we can compute the maximizer $c^{*}$, since
$\begin{gathered}c_{\vec{j}}^{*}:=\operatorname*{arg\,max}_{c}f(w,c)\quad\textrm{s.t.}\quad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\forall m\in\{1,...,M\}\end{gathered}$ (4)
remains a standard convex optimization problem. The final maximum can then be found by aggregation, namely $c^{*}=\operatorname*{arg\,max}_{\vec{j}\in\mathcal{J}}f(w,c_{\vec{j}}^{*})$. While $|\mathcal{J}|=\prod_{k=1}^{K}J_{k}$, we note that most cases of late-stage fusion in practice tend to have a limited number of views, most typically $K=2$ or $K=3$. This coupled with the fact that computing over these indices is trivially parallelizable means this approach is still computationally tractable. The full procedure is outlined in Algorithm 2.
Algorithm 2 Predict-Then-Optimize Under MVCP Prediction Regions
1:Inputs: Context $x$, CGMs $\{q_{k}(C\mid X_{k})\}_{k=1}^{K}$, Optimization steps $T$, Sample counts $\{J_{k}\}_{k=1}^{K}$, MVCP quantile $\widehat{\mathcal{Q}}:=\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$
2:$w\sim U(\mathcal{W}),\{\{\widehat{c}_{jk}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X_{k})\}_{k=1}^{K},\mathcal{J}=\prod_{k=1}^{K}[J_{k}]$
3:for $t\in\{1,\ldots T\}$ do
4:     $\left\{c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}\left\{f(w,c)\mid u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\forall m\in\{1,...,M\}\right\}\right\}_{\vec{j}\in\mathcal{J}}$
5:     $c^{*}\leftarrow\operatorname*{arg\,max}_{c_{\vec{j}}^{*}}f(w,c_{\vec{j}}^{*})$
6:     $w\leftarrow\Pi_{\mathcal{W}}(w-\eta\nabla_{w}f(w,c^{*}))$
7:Return $w$
4 Experiments
We now study MVCP empirically across several tasks, demonstrating its improvement over alternate approaches in a synthetic classification task (Section 4.1), in predict-then-optimize tasks (Section 4.2), and in a classification task over a non-synthetic dataset of interest, namely Handwritten [50] (Section 4.3). To allow for fair comparison, we note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment. This in turn means that some care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for MVCP, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction. For this reason, we fixed the split to be $20\%$-$80\%$ in experiments. Further details are provided in Appendix E, and code is available at: https://github.com/yashpatel5400/fusioncp
4.1 Synthetic Classification Task
We first consider a synthetic classification task, fitting multinomial logistic regression models $\widehat{f}_{k}:\mathcal{X}_{k}\rightarrow\mathcal{Y}$, where $\mathcal{X}=\mathbb{R}^{120}$ and $\mathcal{Y}=\{1,...,20\}$. We consider $K$-view classification problems, with $K\in\{2,...,6\}$, where the views here partition the input dimension into subspaces of dimension $120/K$, namely $\mathcal{X}_{i}=\mathbb{R}^{1+(d/K)(i-1):(d/K)i}$ for $i\in\{1,...,K\}$. We compare MVCP to four baselines: a fusion-free approach that conformalizes the view with the highest predictive accuracy; an approach using a predefined “Norm” fusion algorithm as described in Section 3.1, namely with $g_{1}(s)=(\sum_{k=1}^{K}s_{k}^{2})^{1/2}$; another similarly using a predefined “Sum” fusion algorithm with $g_{2}(s)=\sum_{k=1}^{K}s_{k}$; and one that conformalizes a fused predictor as also discussed in Section 3.1, with $\mathcal{F}=(\widehat{f}_{1}(x)+\dots+\widehat{f}_{K}(x))/K$. We perform conformalization using the standard classification score function across all approaches, namely $s_{k}(x,y)=\sum_{j=1}^{l}\widehat{f}_{k}(x)_{\pi_{j}(x,k)}$ where $y=\pi_{l}(x,k)$ and $\pi(x,k)$ is the permutation of $\{1,\ldots,20\}$ that sorts $\widehat{f}_{k}\left(x\right)$ from most likely to least likely. The results are presented in Table 1.
Table 1: Coverages across tasks for $\alpha=0.10$ are shown in the left table, where coverage was assessed over a batch of 300 i.i.d. test samples. Average prediction set sizes are shown in the right table, averaged over a batch of 300 i.i.d. test samples with standard deviations in parentheses.
$K$ Best View Pre-Fused Sum Norm MVCP
2 0.89 0.91 0.91 0.91 0.90
3 0.89 0.91 0.91 0.91 0.90
4 0.89 0.91 0.90 0.90 0.91
5 0.88 0.90 0.90 0.90 0.90
6 0.89 0.91 0.90 0.90 0.90
Best View Pre-Fused Sum Norm MVCP
10.84 (0.34) 8.11 (0.56) 8.73 (0.42) 8.72 (0.42) 7.58 (0.42)
12.36 (0.30) 8.49 (0.44) 9.05 (0.32) 8.83 (0.35) 7.68 (0.42)
13.32 (0.30) 8.75 (0.51) 9.04 (0.47) 8.41 (0.50) 8.07 (0.50)
13.90 (0.19) 9.01 (0.59) 8.91 (0.63) 8.45 (0.59) 8.37 (0.58)
14.28 (0.27) 9.26 (0.39) 9.04 (0.68) 8.71 (0.71) 8.71 (0.75)
We see that all the approaches exhibit the desired coverage guarantees. However, the multivariate score approaches consistently produce significantly smaller prediction regions than the individually conformalized input views, with MVCP consistently producing the smallest sets on average.
4.2 MVCP Predict-Then-Optimize
We now study the use of MVCP across predict-then-optimize tasks, specifically considering the fractional knapsack problem as in the original conformal predict-then-optimize paper [26]:
$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(X)}\quad-\widehat{c}^{T}w$ (5)
[…]
where $p\in\mathbb{R}^{n}$ and $B>0$. The distributions $\mathcal{P}(C)$ and $\mathcal{P}(X\mid C)$ are again taken to be those from various simulation-based inference (SBI) benchmark tasks provided by [51], chosen as they have $\mathcal{P}(C\mid X)$ with complex structure, fully described in Appendix C. We specifically study the Gaussian Mixture distributional setup across increasing dimensions to highlight the capacity for MVCP to succeed even for complex combinations of distributions.
We then consider $K=2$ views, namely with $\mathcal{X}=\mathbb{R}^{d}$, $\mathcal{X}_{1}=\mathbb{R}^{1:d_{1}}$, and $\mathcal{X}_{2}=\mathbb{R}^{d_{1}+1:d}$. For each $\mathcal{X}_{k}$, we train a generative model $q_{k}(C\mid X_{k})$, specifically a normalizing flow. We then compare four distinct ways of formulating $\mathcal{C}(X)$, namely by conformalizing each view individually, by conformalizing a fused map, by fusing the scores via a sum, and by fusing with MVCP, all using the score given by Equation 3, specifically taking $g_{k}$ to be the $\ell^{2}$ metric. The “Pre-fused” strategy employed here is again a mean aggregation, i.e. a sample $\widehat{c}\sim\mathcal{F}(q_{1},q_{2})$ is drawn via $\widehat{c}_{1}\sim q_{1}(C\mid X_{1}),\widehat{c}_{2}\sim q_{2}(C\mid X_{2})$ and $\widehat{c}$ defined as $(\widehat{c}_{1}+\widehat{c}_{2})/2$.
As discussed in Section 3.3, the prediction regions are not explicitly constructed in regression settings, rendering the reporting of prediction region sizes impossible. We instead compare the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.2. This quantity measures the conservatism of the produced robust optimal value and is bounded in $[0,1]$ in this setting, with the upper bound seen as Equation 5 can trivially be solved by simply taking $w=0$. The results are presented in Table 2.
Table 2: Coverages across tasks for $\alpha=0.05$ are shown in the left table, where coverage was assessed over a batch of 500 i.i.d. test samples. Suboptimality gap proportions ($\Delta_{\%}$) are shown in the right table, with medians reported over a batch of 500 i.i.d. test samples and median average deviations in parentheses.
View 1 View 2 Pre-Fused Sum MVCP
Gaussian Mixture (1) 0.964 0.958 0.952 0.964 0.976
Gaussian Mixture (2) 0.938 0.938 0.93 0.95 0.94
Gaussian Mixture (3) 0.944 0.972 0.936 0.944 0.948
Gaussian Mixture (4) 0.942 0.94 0.914 0.95 0.924
View 1 View 2 Pre-Fused Sum MVCP
0.901 (0.1003) 0.886 (0.1144) 1.0 (0.0) 0.702 (0.1403) 0.542 (0.0946)
0.962 (0.0453) 1.0 (0.0155) 1.005 (0.0019) 0.893 (0.1084) 0.874 (0.0792)
1.011 (0.018) 0.997 (0.0208) 1.011 (0.0035) 0.838 (0.0735) 0.868 (0.047)
1.026 (0.0181) 0.987 (0.0327) 1.018 (0.0066) 1.003 (0.0144) 0.946 (0.0286)
We see that all methods produce the coverage guarantees in accordance to Theorem 3.1. The resulting suboptimality gaps additionally exhibit two notable characteristics. The first is that MVCP produces more informative prediction regions, and hence less conservative robust upper bounds, if the views in consideration are individually informative, that is, each has $\Delta_{\%}<1$. While the sum conformal fusion exhibits a similar pattern, its effective reduction is consistently less than that provided by MVCP. Additionally, in cases where one of the two views is uninformative, i.e. $\Delta_{\%}=1$, MVCP nearly retains the performance of the informative view, as the quantile envelope adapts in considering all projection directions. The sum fusion technique, on the other hand, significantly degrades from the performance of the optimal view.
4.3 Real Data Classification
We now conduct experiments on the Handwritten dataset [50]. This dataset consists of features of 2000 handwritten digits $(0-9)$ (200 examples per class) with 6 different views of each image extracted from a collection of Dutch utility maps, described fully in Appendix D. The baseline approaches parallel those presented in Section 4.1. The results are compiled in Table 3. As in the synthetic classification experiment, we see that considering multiple views produces significantly more informative prediction regions than solely conformalizing a single view, especially when done in the manner prescribed by MVCP.
Table 3: Coverages across tasks for $\alpha=0.10$ are shown in the left table, where coverage was assessed over a batch of 295 i.i.d. test samples. Average prediction set sizes are shown in the right table, averaged over a batch of 295 i.i.d. test samples with standard deviations in parentheses.
$K$ Best View Pre-Fused Sum Norm MVCP
2 0.90 0.91 0.91 0.91 0.90
3 0.90 0.90 0.90 0.90 0.90
4 0.91 0.90 0.90 0.90 0.90
5 0.91 0.91 0.91 0.90 0.90
6 0.90 0.91 0.90 0.90 0.91
Best View Pre-Fused Sum Norm MVCP
3.39 (0.09) 3.69 (0.11) 3.52 (0.13) 3.51 (0.13) 2.68 (0.17)
3.39 (0.09) 3.48 (0.09) 2.97 (0.09) 2.95 (0.09) 1.84 (0.10)
3.28 (0.14) 3.42 (0.05) 2.87 (0.05) 2.83 (0.05) 1.64 (0.12)
3.28 (0.14) 3.12 (0.07) 2.73 (0.08) 2.68 (0.07) 1.55 (0.06)
3.19 (0.17) 3.07 (0.11) 2.71 (0.06) 2.64 (0.05) 1.60 (0.07)
5 Discussion
We have presented MVCP, a framework for producing informative prediction regions in late-stage fusion pipelines. We additionally demonstrated the generality of the proposed pipeline in settings of both classification and robust predict-then-optimize regression tasks. This work suggests many directions for extension. One point of interest is the integration of MVCP with other downstream regression applications. For instance, [52] proposed an end-to-end differentiable extension to [26]: extending MVCP for integration to their framework would be of great interest. Additionally, given the prevalence of late-stage fusion in robotics applications, an exciting applied avenue of research would investigate the optimization of MVCP to leverage it in settings of robust control.
6 Acknowledgements
We acknowledge the support of the College of Literature, Science, and the Arts at the University of Michigan via the Meet the Moment Research Initiative.
Appendix A Multivariate Visualization
We walk through a visual presentation of the approach below to supplement the textual description in the main text. As discussed, we start with a collection of multivariate calibration scores $\mathcal{S}_{C}$ with $s\in\mathcal{S}$ being $\in\mathbb{R}^{K}$. For the purposes of visualization in this section, we have $K=2$. We first partition the score evaluations $\mathcal{S_{C}}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$ into a subset used to define the pre-ordering and the remainder to define the multivariate quantile.
Figure 2: The calibration score evaluations are first split between those used to define the pre-ordering (green) and those used to define the final multivariate quantile (red).
We first wish to define the pre-ordering over $\mathcal{S}_{C}^{(1)}$. As described in the main text, the goal is to define this using an indexed family of sets $\mathcal{A}_{t}$ with index $t\in\mathbb{R}$, after which the multivariate quantile reduces to the univariate quantile formulation. To ensure the final envelope over $\mathcal{S}_{C}^{(2)}$ remains as tight as possible, we wish to define this family in a data-driven fashion. Critically, the shape of this tightest envelope around $\mathcal{S}_{C}^{(2)}$ will vary across $\alpha$, meaning we must define the family separately for each choice of $\alpha$. We expect the contour of the tightest $\alpha$ envelope for $\mathcal{S}_{C}^{(1)}$ will be similar to that over $\mathcal{S}_{C}^{(2)}$, motivating such a choice to define the indexing family. To do this, we project $\mathcal{S}_{C}^{(1)}$ along a number of directions, finding the $\beta$ quantile along each, in turn defining a half-plane, where $\beta$ is as described in Section 3.2.
Figure 3: The pre-ordering points are projected across a number of directions, after which the $\beta$ quantile is used to define a direction quantile. This defines a half-plane of points that are in the region (blue) and those outside (red).
We then iteratively update $\beta$ in the manner described in Algorithm 1 to obtain $\beta^{*}$, namely the minimum value for which the region given by the intersection of the corresponding half-planes covers roughly $1-\alpha$ of $\mathcal{S}_{C}^{(1)}$.
Figure 4: We use the intersection of hyperplanes to define the quantile envelope, seeking $\beta^{*}$ that achieves the desired coverage.
Once this $1-\alpha$ quantile envelope of $\mathcal{S}_{C}^{(1)}$ is found, we define $\mathcal{A}_{1}$ to be such an envelope, with which future points can now be partially ordered. That is, for any point $s\in\mathbb{R}^{K}$ notice that we can unambiguously associate it with $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$. Intuitively, this is the $t$ where the contour “intersects” $s$. Notably, now that the partial ordering has been defined, the points of $\mathcal{S}_{C}^{(1)}$ are no longer used. It would be of interest to investigate whether a concurrent definition of the partial ordering and final calibration is possible without such data splitting in future work.
Figure 5: Using the quantile envelope, the family of nested sets $\mathcal{A}_{t}$ is defined, in turn defining a partial ordering over $\mathbb{R}^{K}$.
With this $\mathcal{A}_{t}$, we find the final $\widehat{q}$ simply by mapping the points of $\mathcal{S}_{C}^{(2)}$ to their corresponding $t(s)$ values in the aforementioned fashion and performing standard conformal prediction. As discussed, if the envelope has a similar structure to that found over $\mathcal{S}_{C}^{(1)}$, the envelope should be adjusted by only a minor amount.
Figure 6: Using the nested family of sets, we expand or contract the envelope appropriately using the data of $\mathcal{S}_{C}^{(2)}$ to find the final adjustment factor.
Appendix B Multivariate Score Coverage
The proof of the multivariate extension of conformal prediction follows in precisely the same manner as that of standard conformal prediction with the pre-order $\lesssim$ replacing the complete ordering used in traditional conformal prediction.
Theorem B.1.
Suppose $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$ are exchangeable. Assume further that $K$ maps $s_{j}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined. Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\}_{(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}}$. For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$. Then, denoting $\mathcal{C}(X):=\{y:s(X,y)\in\mathcal{Q}(\alpha)\}$,
$\mathcal{P}_{X,Y}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha.$ (6)
Proof.
Denote $s_{i}=s(X_{i},Y_{i})$ for each $i=1,...,n$ and $s^{\prime}=s(X^{\prime},Y^{\prime})$. We define $t_{i}=\inf\{t\geq 0:s_{i}\in\mathcal{A}_{t}\}$ and $t^{\prime}=\inf\{t\geq 0:s^{\prime}\in\mathcal{A}_{t}\}$. We consider the case that $\mathcal{P}_{X,Y}(t_{i}\neq t_{j})=1$, that is, that the probability of ties is a probability measure 0 set. Without loss of generality, we then assume the scores are sorted according to the assumed pre-order, namely that $s_{1}\lesssim s_{2}\lesssim...\lesssim s_{N_{\mathcal{C}}}$, or equivalently $t_{1}\leq t_{2}\leq...\leq t_{N_{\mathcal{C}}}$. We then again have that $\widehat{t}=t_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}$ if $\alpha>1/(N_{\mathcal{C}}+1)$ and $\widehat{t}=\infty$ otherwise. In the latter case, coverage is trivially satisfied. In the former case, we see
$\mathcal{P}_{X,Y}(Y^{\prime}\in\mathcal{C}(X^{\prime}))=\mathcal{P}_{X,Y}(s(X,y)\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil})=\mathcal{P}_{X,Y}(t^{\prime}\leq t_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}).$ (7)
By the assumed exchangeability of $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$, we have that
$\mathcal{P}_{X,Y}(t^{\prime}\leq t_{k})=\frac{k}{N_{\mathcal{C}}+1},$ (8)
for any $k$. From here, we have the desired conclusion that
$\mathcal{P}_{X,Y}(s(X,y)\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil})=\left(\frac{1}{N_{\mathcal{C}}+1}\right)(\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil)\geq 1-\alpha,$ (9)
completing the proof as desired. ∎
Appendix C Simulation-Based Inference Benchmarks
The benchmark tasks are a subset of those provided by [53]. For convenience, we provide brief descriptions of the tasks curated by this library; however, a more comprehensive description of these tasks can be found in their manuscript.
C.1 Gaussian Mixture
A mixture of two Gaussians, with one having a much broader covariance structure:
$\text{{Prior}: }\beta\sim\mathcal{U}(-10,10)$
$\text{{Simulator}: }x\mid w\sim 0.5\mathcal{N}(x\mid w,I)+0.5\mathcal{N}(x\mid w,.01I)$
For any Gaussian Mixture ($k$) distribution referenced in Section 4, we combined separate instances of this setup, where the typical setup has $\beta\in\mathbb{R}^{2}$. In other words, we generate $x:=[x_{1},...,x_{k}]$ by drawing a prior over a similarly partitioned $\theta:=[\theta_{1},...,\theta_{k}]$ and taking $x_{i}\sim\mathcal{P}(X_{i}\mid\theta_{i})$ according to the above distribution.
Appendix D Real Data Featurization
The data features of the Handwritten dataset [50] are as follow:
  • 47 Zernike moments
  • 76 Fourier coefficients of the character shapes
  • 6 morphological features
  • 240 pixel averages in $2\times 3$ windows
  • 64 Karhunen-Love coefficients
  • 216 profile correlations
Appendix E Training Details
All encoders were implemented in PyTorch [54] with a Neural Spline Flow architecture. The NSF was built using code from [55]. Specific architecture hyperparameter choices were taken to be the defaults from [55] and are available in the code. Optimization was done using Adam [56] with a learning rate of $10^{-3}$ over 5,000 training steps. Minibatches were drawn from the corresponding prior $\mathcal{P}(Y)$ and simulator $\mathcal{P}(X\mid Y)$ as specified per task in the preceding section. Training these models required between 10 minutes and two hours using an Nvidia RTX 2080 Ti GPUs for each of the SBI tasks.

### v2

@@
-Conformalized Late Fusion Multi-View Learning
+Conformal Prediction for Ensembles: Improving Efficiency via Score-Based Aggregation
 Abstract
-Uncertainty quantification for multi-view learning is motivated by the increasing use of multi-view data in scientific problems.
-A common variant of multi-view learning is late fusion: train separate predictors on individual views and combine them after single-view predictions are available.
-Existing methods for uncertainty quantification for late fusion often rely on undesirable distributional assumptions for validity.
+Distribution-free uncertainty estimation for ensemble methods is increasingly desirable due to the widening deployment of multi-modal black-box predictive models.
 Conformal prediction is one approach that avoids such distributional assumptions.
-However, naively applying conformal prediction to late-stage fusion pipelines often produces overly conservative and uninformative prediction regions, limiting its downstream utility.
-We propose a novel methodology, Multi-View Conformal Prediction (MVCP), where conformal prediction is instead performed separately on the single-view predictors and only fused subsequently.
-Our framework extends the standard scalar formulation of a score function to a multivariate score that produces more efficient downstream prediction regions in both classification and regression settings.
-We then demonstrate that such improvements can be realized in methods built atop conformalized regressors, specifically in robust predict-then-optimize pipelines.
+Methods for conformal aggregation have in turn been proposed for ensembled prediction, where the prediction regions of individual models are merged as to retain coverage guarantees while minimizing conservatism.
+Merging the prediction regions directly, however, sacrifices structures present in the conformal scores that can further reduce conservatism.
+We, therefore, propose a novel framework that extends the standard scalar formulation of a score function to a multivariate score that produces more efficient prediction regions.
+We then demonstrate that such a framework can be efficiently leveraged in both classification and predict-then-optimize regression settings downstream and empirically show the advantage over alternate conformal aggregation methods.
 1 Introduction
-Multi-view machine learning is becoming increasingly common with the advent of multi-modal language models and conditional generative models [1, 2, 3, 4, 5].
-In addition, such techniques are being increasingly used in more safety-critical settings, such as medicine [6, 7, 8, 9] and robotics [10, 11, 12, 13].
-These use cases motivate the development of uncertainty quantification methods for multi-view learning.
-The most common division of multi-view learning is along the axis of when fusion of the multiple views is performed.
-We can either learn a fused data representation or we can fuse the predictions from the separate views.
-The latter approach is referred to as “late fusion”.
-It is desirable in limited resource settings, where the separately trained models can be combined in a light-weight fashion for downstream tasks [14, 15, 16, 17].
-Increasing interest is now being placed on quantifying uncertainty for such models [18, 19, 20, 21, 22].
-Such methods, however, rely on distributional assumptions to provide coverage guarantees or are computationally intractable, limiting their practical utility.
-These limitations, therefore, demand a more universal, yet informative, uncertainty quantification method.
-One such method for performing distribution-free uncertainty quantification is conformal prediction, which provides a principled framework for producing distribution-free prediction regions with marginal frequentist coverage guarantees [23, 24].
-By using conformal prediction on a user-defined score function $s(x,y)$ and obtaining an empirical $1-\alpha$ quantile $\widehat{q}(\alpha)$ of $s(x,y)$ over a calibration set $\mathcal{D}_{C}$, prediction regions $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}(\alpha)\}$ attain marginal coverage guarantees.
+Ensemble methods are an oft-used class of statistical modeling techniques due to their ability to reduce variance or improve predictive accuracy [1, 2, 3].
+Such methods are increasingly being coupled with complex, black-box models, such as in multi-modal language models [4, 5, 6, 7, 8].
+Couplings of this sort are seeing ever-widening deployment in safety-critical settings, such as medicine [9, 10, 11] and robotics [12, 13, 14].
+Increasing interest is, therefore, now being placed on quantifying uncertainty for such models [15, 16, 17, 18, 19].
+Towards this end, methods of uncertainty quantification have arisen, such as deep ensembles and committee estimation [20, 21, 22].
+Such methods, however, sacrifice generality with the imposition of distributional assumptions, motivating the need for distribution-free uncertainty quantification for ensemble methods.
+One method for performing distribution-free uncertainty quantification is conformal prediction, which provides a principled framework for producing distribution-free prediction regions with marginal frequentist coverage guarantees [23, 24].
+By using conformal prediction on a user-defined score function, prediction regions attain marginal coverage guarantees.
 While calibration is guaranteed from this procedure, predictive efficiency, quantified as the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
-We extend the conformal prediction line of work to enable its practical use in late-stage multi-view prediction problems.
-In particular, we propose a multivariate extension to the score function notion introduced in traditional conformal prediction, with a corresponding extension of quantiles to “quantile envelopes.” We demonstrate that this formulation retains the desired distribution-free coverage guarantees typical of standard conformal prediction and that the resulting prediction regions can be used efficiently in both classification and regression settings.
-Our contributions are:
-Figure 1: MVCP provides a principled extension to the standard conformal prediction pipeline by leveraging ideas from higher-dimensional quantile regression to define quantile envelopes $\widehat{\mathcal{Q}}$ instead of purely scalar quantiles $\widehat{q}$.
+As a result, methods have arisen to perform conformal model aggregation, which both provide uncertainty estimates of the ensembled predictions and do so in ways as to minimize the prediction region size [25, 26, 27, 28, 29].
+While such approaches succeed in reducing the prediction region size over naive aggregation, they all aggregate the separately conformalized prediction regions of the predictors in the ensemble.
+In doing so, they forgo the possibility of automatically leveraging shared structure amongst the scores of the individual predictors, resulting in overly conservative prediction regions.
+Figure 1: CSA provides a principled extension to the standard conformal prediction pipeline by leveraging ideas from higher-dimensional quantile regression to define quantile envelopes $\widehat{\mathcal{Q}}$ instead of scalar quantiles $\widehat{q}$.
 It does so by evaluating a collection of score functions (here $s_{1}$ and $s_{2}$) over the calibration dataset to define $\mathcal{S}$, finding quantiles $\{\widehat{q}_{m}\}$ over a set of projection directions $\{u_{m}\}$, and taking $\widehat{\mathcal{Q}}$ to be the intersection of the resulting half-planes $H(u_{m},\widehat{q}_{m})$.
 These quantile envelopes result in more informative prediction regions that can be used in downstream tasks.
-  • Providing a multivariate extension to conformal prediction score functions (MVCP) that retains coverage guarantees and can be applied to late-stage multi-view prediction problems.
-  • Demonstrating how the resulting prediction regions can be efficiently leveraged for downstream classification and regression tasks.
-  • Demonstrating the empirical performance of the MVCP framework across a suite of tasks.
+We instead propose to perform aggregation in score space by extending traditional conformal prediction to consider a multivariate score function and defining prediction regions using “quantile envelopes” in place of scalar quantiles.
+Doing so enables efficient, data-driven, automated conformal model aggregation.
+We demonstrate that this formulation retains the desired distribution-free coverage guarantees typical of standard conformal prediction and that the resulting prediction regions can be used efficiently in both classification and regression settings.
+Our contributions, thus, are:
+  • Providing a multivariate extension to conformal prediction, dubbed “conformal score aggregation” (CSA), that leverages quantile envelopes to enable data-driven, informative uncertainty estimation for model ensembles while retaining coverage guarantees.
+  • Demonstrating how the prediction regions resulting from CSA can be efficiently leveraged in downstream predict-then-optimize regression tasks.
+  • Demonstrating the empirical improvement of the CSA framework over alternate conformal aggregation strategies across classification and regression settings.
 2 Background
 2.1 Conformal Prediction
-Coverage guarantees of uncertainty quantification methods generally rely on distributional assumptions, often via asymptotics or explicit model specification.
+Coverage guarantees of uncertainty quantification methods generally rely on distributional assumptions, often via asymptotics or explicit specification.
 To alleviate the need for such restrictive assumptions, interest in finite-sample, distribution-free uncertainty quantification methods has risen.
@@
 Conformal prediction serves as a wrapper around such predictors, producing prediction regions $\mathcal{C}(x)$ that have formal guarantees of the form $\mathcal{P}_{X,Y}(Y\notin\mathcal{C}(X))\leq\alpha$ for some prespecified level $\alpha$.
-To achieve this, “split conformal” partitions the dataset $\mathcal{D}$ into a training set $\mathcal{D}_{T}$ and a calibration set $\mathcal{D}_{C}$.
-The former plays the role traditionally played by $\mathcal{D}$, namely serving as data used to fit $\widehat{f}$.
-The latter is used much in the way a validation set is typically employed in practice.
-Formally, users of conformal prediction must design a “score function” $s(x,y)$, which should quantify “test error,” often in a domain-specific manner.
+To achieve this, “split conformal” partitions the dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ into a training set $\mathcal{D}_{T}$ and a calibration set $\mathcal{D}_{C}$.
+The former serves as the data used to fit $\widehat{f}$.
+Users of conformal prediction must then design a “score function” $s(x,y)$, which should quantify “test error”, often in a domain-specific manner.
 For instance, a simple score function for a regression setting would be $s(x,y)=\|\hat{f}(x)-y\|$.
-This score function is then evaluated across the calibration set to define $\mathcal{S}=\{s(x,y)\mid(x,y)\in\mathcal{D}_{C}\}$.
-For a desired coverage of $1-\alpha$, we then take $\widehat{q}$ to be the $\lceil(|\mathcal{D}_{C}|+1)(1-\alpha)\rceil/|\mathcal{D}_{C}|$ quantile of $\mathcal{S}$, with which prediction regions for future test queries $x$ can be defined as $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}\}$.
-Under the exchangeability of the score of a test point $s(x^{\prime},y^{\prime})$ with $\mathcal{S}$, we have the desired finite-sample probabilistic guarantee that $1-\alpha\leq\mathcal{P}_{X,Y}(Y\in\mathcal{C}(X))$.
-While this guarantee holds for any $s(x,y)$, the informativeness of the resulting prediction regions, formally defined as the inverse expected Lebesgue measure across $X$, i.e.
+This score function is then evaluated across the calibration set to define $\mathcal{S}_{C}=\{s(x,y)\mid(x,y)\in\mathcal{D}_{C}\}$.
+For a desired coverage of $1-\alpha$, we then take $\widehat{q}$ to be the $\lceil(|\mathcal{D}_{C}|+1)(1-\alpha)\rceil/|\mathcal{D}_{C}|$ quantile of $\mathcal{S}_{C}$, with which prediction regions for future test queries $x$ can be defined as $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}\}$.
+Under the exchangeability of the score of a test point $s(X^{\prime},Y^{\prime})$ with $\mathcal{S}_{C}$, we have the desired finite-sample probabilistic guarantee that $1-\alpha\leq\mathcal{P}_{X^{\prime},Y^{\prime}}(Y^{\prime}\in\mathcal{C}(X^{\prime}))$.
+While this guarantee holds for any $s(x,y)$, the informativeness of the resulting prediction regions, quantified as the inverse expected Lebesgue measure across $X$, i.e.
 $\left(\mathbb{E}[\mathcal{L}(\mathcal{C}(X))]\right)^{-1}$, is intimately tied to its specification [24].
-Thus, much of the practical challenge of conformal prediction relates to choosing a score function that retains coverage while minimizing prediction region size.
+Thus, much of the challenge of conformal prediction relates to choosing a score function that retains coverage while minimizing region size.
+2.2 Quantile Envelopes
+Generalizations of quantiles have a long history in statistics [30, 31].
+Unlike univariate data, multivariate data does not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering in higher dimensional spaces.
+To account for this, one class of multivariate quantiles, known as “direction quantiles,” prescribes directions along which data points can be compared [32, 33, 34].
+That is, points $x\in\mathbb{R}^{n}$ are projected along some direction $u\in\mathcal{S}^{n-1}$, after which the standard notion of a quantile can be leveraged.
+The notion of the $\alpha$ quantile for a random variable $X\in\mathcal{X}$ clearly then depends on this choice of direction $u$, which we denote by $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$.
+When there is no ambiguity, we just denote it as $Q(\alpha,u)$.
+In cases where domain knowledge assigns an unambiguous notion of multivariate ordering, such knowledge can be encoded into the choice of $u$.
+In other cases, however, a more agnostic approach is employed: rather than choosing any particular direction, all directions are considered simultaneously.
+For any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$.
+To consider all directions simultaneously, the quantile envelope is then defined to be
+$D(\alpha)=\bigcap_{u\in\mathcal{S}^{n-1}}H(u,Q(\alpha,u)).$ (1)
+Notably, while each individual $H(u,Q(\alpha,u))$ captures $1-\alpha$ of the points, $D(\alpha)$ does not, as it is the intersection thereof and hence captures $<1-\alpha$ of the mass.
+If $1-\alpha$ combined coverage is sought, a correction, such as Bonferroni adjustment, is used for the individual planes.
 2.3 Predict-Then-Optimize
 Prediction regions produced from conformalized predictors are often leveraged differently across classification and regression settings.
-In the case of classification, “regions” simply constitute a subset of the label space, making its direct use by end users straightforward, as demonstrated in works such as [25].
+In the case of classification, “regions” simply constitute a subset of the label space, making their direct use by end users straightforward, as demonstrated in works such as [35].
 In high-dimensional regression settings, however, prediction regions become harder to use directly; for this reason, recent works have started shifting focus to using them in their implicit forms.
@@
 An example of such a setting is if an optimal labor allocation $w^{*}$ is sought based on predicted demand $\widehat{c}$ from observed transactions $x$ in a food delivery platform.
-While the predicted $\widehat{c}$ is often simply trusted in the downstream optimization, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to highly suboptimal decision-making.
+While the predicted $\widehat{c}$ is often simply trusted in the downstream optimization, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to suboptimal decision-making.
 For this reason, recent interest has been placed on studying a “robust” alternative to this formulation [37, 38, 39].
-Following this line of work, [26] proposed studying the following formulation: $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with the uncertainty region $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$.
+Following this line of work, [36] proposed studying $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$.
 The primary consequence of doing so was having a probabilistic guarantee on the suboptimality gap.
 That is, denoting the suboptimality gap by $\Delta(x,c):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})-\min_{w}f(w,c)$, where $c$ is the true parameter corresponding to $x$, they established that leveraging conformal uncertainty regions guarantees $\mathcal{P}_{X,C}\left(0\leq\Delta(X,C)\leq L\mathrm{\ diam}(\mathcal{C}(X))\right)\geq 1-\alpha$ if $f(w,c)$ is an objective function that is convex-concave and $L$-Lipschitz in $c$ for any fixed $w$.
-2.3 Multi-View Learning
-We provide a brief overview of multi-view learning herein, deferring to recent surveys of the field for a more comprehensive review thereof [30].
-In general, the focus of multi-view machine learning is to solve prediction problems when different, often non-overlapping, representations of the same underlying “object” are observed.
-Methods in this space bifurcate between those defined together with some downstream supervised task and those seeking to find an abstract “fused” representation that lends itself naturally for arbitrary downstream use cases.
-We focus herein on the former category.
-For instance, one may wish to classify a captioned image, where the image and caption represent two distinct “views” of the same underlying abstract object in this case.
-Formally, therefore, we consider the multi-view prediction problem as seeking a map $\mathcal{X}\rightarrow\mathcal{Y}$ with $\mathcal{X}=\mathcal{X}_{1}\cup...\cup\mathcal{X}_{K}$.
-As previously mentioned, fusion approaches are generally either early-stage or late-stage.
-In the former, we seek to learn some common, often abstract, space into which these separate views can be mapped for downstream prediction, namely seeking a map $g_{\phi}$ that maps $g_{\phi}:\mathcal{X}_{1}\times...\times\mathcal{X}_{K}\rightarrow\mathcal{Z}$.
-Such a fusion map is then assumed to be usable for the downstream prediction task, namely through composition with $f_{\theta}:\mathcal{Z}\rightarrow\mathcal{Y}$ such that the overall prediction model can be defined as $f\circ g$.
-Traditionally, $g$ was hand-crafted, leveraging domain-specific insights.
-However, increasingly, $f$ and $g$ are simply subsets of a neural network trained end-to-end via gradient descent, such as in [31, 32, 33, 34, 35, 36].
-As discussed previously, this fusion approach is not universally desirable in all circumstances though, such as when models for individual views have already been curated after computationally intensive training and hyperparameter tuning [14, 15, 16, 17].
-In such cases, late-stage fusion approaches can be preferred, where separate predictors $f_{i}:\mathcal{X}_{i}\rightarrow\mathcal{Y}$ have been trained and are fused via a post-hoc algorithm $\mathcal{F}:\mathcal{Y}\times...\times\mathcal{Y}\rightarrow\mathcal{Y}$.
-Approaches to this include decision averaging [37, 38], decision weighting [18, 39, 40], and decision voting [41].
-We focus herein on such late-stage fusion.
 2.4 Related Works
-Work in the conformal prediction literature has yet to leverage the generalized notions of quantiles considered herein.
-Generalizations of quantiles, while introduced first in the context of conformal prediction herein, have a long history in traditional statistics, with recent surveys on approaches available in [42, 43].
-We briefly summarize the relevant pieces of this literature below.
-Unlike univariate data, multivariate data does not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering of points in higher dimensional spaces.
-To account for this, one class of multivariate quantiles, known as “direction quantiles,” prescribes directions along which data points can be compared [44, 45, 46].
-That is, for some direction $u\in\mathcal{S}^{n-1}$, points $x\in\mathbb{R}^{n}$ are projected, after which the standard notion of a quantile can be leveraged.
-The notion of the $\alpha$ quantile for a random variable $X\in\mathcal{X}$ clearly then depends on this choice of direction $u$, which we denote by $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$.
-When there is no ambiguity, we just denote it as $Q(\alpha,u)$.
-In cases where domain knowledge assigns an unambiguous notion of multivariate ordering, such knowledge can be encoded into the choice of $u$.
-In other cases, however, a more agnostic approach is typically employed: rather than choosing any particular direction, all directions are considered simultaneously.
-Formally, for any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$.
-To extend this to consider all directions simultaneously, we then define the quantile envelope to be
-$D(\alpha)=\bigcap_{u\in\mathcal{S}^{n-1}}H(u,Q(\alpha,u)).$ (1)
-To the best of our knowledge, the only work leveraging conformal prediction in the context of multivariate quantiles is [47].
-This work, however, focuses on a fundamentally different problem than that considered herein.
-In particular, they focused on leveraging conformal prediction as a correction step to the output of directional quantile regression with the intended use case being in constructing calibrated prediction regions for multivariate regression problems.
-This work, therefore, was a natural extension to the now classical application of conformal prediction for the correction of univariate quantile regression.
-Our proposed method, however, does the converse: it leverages directional quantiles within the score space to extend conformal prediction itself.
+Ensemble methods consist of $K$ predictors $f_{k}:\mathcal{X}_{k}\rightarrow\mathcal{Y}$; notably, such predictors need not map from the same set of covariates.
+A naive approach for uncertainty quantification would then be to conformalize the ensembled predictor.
+That is, for an ensembling algorithm $\mathcal{F}:\mathcal{Y}^{K}\rightarrow\mathcal{Y}$, a score function $s(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y)$ would be defined.
+Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ would then be marginally calibrated.
+Such an approach, however, lacks some desirable properties.
+In particular, prediction regions $\mathcal{C}(x)$ should have the quality that, if a particular predictor has less uncertainty in its predictions, as is frequently true of ensemble settings where the predictors span multiple input data modalities, upon routing to that predictor, the corresponding size of the prediction region should be smaller than if it had been routed to a different predictor.
+While the naive approach does, in principle, support this property, it ultimately relies on defining an uncertainty-aware ensembling algorithm $\mathcal{F}$.
+In its typical form, however, $\mathcal{F}$ simply takes point predictions $f_{1}(x),...,f_{K}(x)$ in as input, meaning any uncertainty-awareness would need to be baked in a priori into the definition of $\mathcal{F}$ through domain knowledge of the uncertainties of the predictors $f_{1},...,f_{K}$, which can seldom be specified precisely, sacrificing the predictive efficiency of $\mathcal{C}(x)$.
+Conformal model aggregation, thus, seeks to mitigate these deficiencies by aggregating the prediction regions $\mathcal{C}_{1}(x),...,\mathcal{C}_{K}(x)$ rather than the individual point predictions [25, 27, 28, 29].
+While there are several methods in this vein, they can be categorized into one of two general approaches.
+The first line of work seeks to perform model selection, in which a single conformal predictor is selected $\mathcal{C}_{k^{*}}$, typically based on the criterion of minimizing region size $k^{*}:=\operatorname*{arg\,min}_{k}\mathbb{E}[\mathcal{L}(\mathcal{C}_{k}(X))]$ [27, 28].
+Generally, however, methods leveraging the full collection of predictors produce less conservative regions [25, 29].
+Such works aggregate the individual prediction regions into a final region by defining
+$\mathcal{C}(x):=\left\{y\mid\sum_{k=1}^{K}w_{k}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]\geq\widehat{a}\right\}$ (2)
+for weights $\{w_{k}\}\in[0,1]$ such that $\sum_{k=1}^{K}w_{k}=1$ and a threshold $\widehat{a}$.
+Methods then differ in the procedure by which $\{w_{k}\}$ and $\widehat{a}$ are prescribed, several of which were prescribed in [29], whose detailed presentation is deferred to Appendix F for space reasons.
+We note that the methods of [25] are designed for a different setting than that considered herein, namely that in which conformal coverage is sought adaptively over online data streams.
+In this vein, [40] is a recent work that has similarly proposed a vector-score extension as that discussed herein, in which candidate weight vectors $\{w_{m}\}\in\mathbb{R}^{K}$ are searched over for score aggregation.
+That is, a vector $s(x):=(s_{1}(x,y),...,s_{K}(x,y))\in\mathbb{R}^{K}$ of scores $s_{k}(x,y)$ corresponding to each predictor $f_{k}(x)$ is predicted and its aggregate prediction region defined on the projection $\langle w_{m^{*}},s\rangle$ for $w_{m^{*}}$ the weight resulting in the smallest prediction region.
+This method, however, has two shortcomings addressed herein.
+The first is that their method can only be applied in classification settings, whereas our method can be leveraged across both regression and classification problems.
+The second is that their approach only uses a single weighted projection in the end, resulting in suboptimal aggregation and, therefore, conservative prediction regions.
 3 Method
-We now introduce a procedure for uncertainty quantification for late-stage fusion.
-We discuss the naive approach in Section 3.1 and its extension along with the associated coverage guarantees in Section 3.2.
-We then discuss the downstream use of the resulting prediction regions in Section 3.3.
-3.1 Naive Approach
-We now consider the setting described in the background section, where $K$ predictors $f_{1},...,f_{K}$ are trained on the separate views with a late fusion algorithm $\mathcal{F}$, which we assume to be fixed throughout this presentation.
-Note that the validity of the results presented below is independent of the predictive quality of such $f_{i}$.
-The efficiency, however, is highly dependent on it.
-In the standard application of conformal prediction, a score function $s(\mathcal{F}(x),y)$ is defined.
-For instance, if we assume $\mathcal{Y}$ is some metric space with metric $d$, as in a regression task, a naive choice of score function would be
-$s(x,y)=d(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y).$ (2)
-Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ is then marginally calibrated.
-Ideally, however, prediction regions $\mathcal{C}(x)$ should have the quality that, if a particular predictor has less uncertainty in its predictions, as is frequently true of settings leveraging late-stage fusion, upon routing to that predictor, the corresponding size of the prediction region should be smaller than if it had been routed to a different predictor.
-While the framing of Equation 2 does, in principle, support this property, it ultimately relies on defining an uncertainty-aware late fusion algorithm $\mathcal{F}$.
-In its typical form, however, $\mathcal{F}$ simply takes point predictions $f_{1}(x),...,f_{K}(x)$ in as input, meaning any uncertainty-awareness would need to be baked in a priori through domain knowledge of the uncertainties of the predictors $f_{1},...,f_{K}$, which can seldom be specified precisely, in turn sacrificing the downstream predictive efficiency of $\mathcal{C}(x)$.
-For this reason, we seek to produce a data-driven approach of obtaining the uncertainties of the predictors by instead fusing the conformalized predictors.
-In this case, we first define a score function $s_{i}$ for each predictor $f_{i}$.
-For any $x$, therefore, we now have a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$.
+We now introduce a procedure for conformal score aggregation for ensembles.
+We discuss the approach and its coverage guarantees in Section 3.1.
+We then discuss the downstream use of its prediction regions in Section 3.2.
+3.1 Multivariate Score Quantile
+We wish to consider the setting typical of conformal model aggregation, as discussed in Section 2.4, in which predictors $f_{1}(x),...,f_{K}(x)$ and corresponding scores $s_{1}(x,y),...,s_{K}(x,y)$ are defined.
+We further assume a similar premise as that taken by [40], in which the individual scores are stacked into a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$.
 A naive approach would then involve leveraging standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$.
-As in the case of conformalizing the fused predictor directly, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties.
+Similar to the naive conformalization of an ensembled predictor discussed in Section 2.4, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties.
 We instead wish to provide a data-adaptive pipeline to automatically produce such a $g$.
-Formally, denoting by $\mathcal{C}(g):\mathcal{X}\rightarrow\mathcal{Y}$ the conformalized predictor resulting from calibrating with the score fusion function $g$ in some function class $\mathcal{G}$, the goal can be formulated as seeking $g^{*}:=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{X}\left[\mathcal{L}\left(\mathcal{C}\left(g\right)\left(X\right)\right)\right]$.
-There is, however, no approach to solving this abstract formulation, since even the mere evaluation of $\mathbb{E}_{X}[\mathcal{L}(\mathcal{C}(g)(X))]$ for a fixed $g$ is intractable.
-3.2 MVCP: Score Quantile Envelope
-Since finding $g^{*}$ exactly is intractable, we propose a data-adaptive approach that leverages the notion of quantile envelopes from Section 2.4 over the scores $s\in\mathbb{R}^{K}$.
-Importantly, for the remainder of this discussion, we assume the score functions are non-negative, i.e., $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which is typically the case as the score serves as a generalization of the classical idea of a residual.
-Intuitively, the coverage guarantee in this multivariate extension requires us to define a pre-ordering $\lesssim$ over $\mathbb{R}^{K}$, where we only lose the antisymmetric axiom of a total ordering.
-Upon defining such a pre-ordering, the coverage guarantee follows in using the appropriate quantile under this modified sorting.
-Crucially, however, the problem of choosing this pre-ordering closely parallels that of choosing $g^{*}$, where a poorly chosen pre-ordering will result in overly conservative quantile envelopes and, hence, prediction regions.
-For instance, using a lexicographical ordering $\lesssim_{\mathrm{Lex}}$ will result in axis-aligned hyper-rectangular quantile regions and similarly a norm-based ordering $\lesssim_{\mathrm{Norm}}$, defined as $s_{1}\lesssim_{\mathrm{Norm}}s_{2}\iff||s_{1}||_{2}\leq||s_{2}||_{2}$, in hyperspherical regions.
-Rather than explicitly prescribing a pre-ordering, $\lesssim$ can be implicitly defined by prescribing an indexed family of nested sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, such that $\mathcal{A}_{t_{1}}\subset\mathcal{A}_{t_{2}}$ for $t_{1}\leq t_{2}$.
-We then define $\lesssim$ by stating $s_{1}\lesssim s_{2}$ if $\forall t$, $s_{2}\in\mathcal{A}_{t}\implies s_{1}\in\mathcal{A}_{t}$.
-While not strictly necessary, we restrict ourselves to the case where $\mathcal{A}_{t}$ is a family of sets defined by a fixed “base contour” that scales outward with $t$.
-Under this restriction, performing a comparison of $s_{1},s_{2}\in\mathbb{R}^{K}$ in the aforementioned fashion, i.e.
-checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the $t$ for which $\mathcal{A}_{t}$ “intersects” $s$.
+Importantly, we heretofore assume the score functions are non-negative, i.e., $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which is typically the case as the score serves as a generalization of the residual.
+We highlight that many of the details of the method presented below are geometric in nature and are more easily understood with the supplement of diagrams.
+We have, thus, provided an accompanying visual walkthrough of the procedure in Appendix A to clarify its presentation.
+3.1.1 Score Partial Ordering
+Intuitively, our method seeks to directly generalize the approach of split conformal, by “ordering” the collection of multivariate calibration scores and taking the $1-\alpha$ score under such an ordering to be a threshold $\widehat{\mathcal{Q}}$ with which prediction regions are then implicitly defined.
+Formally, the multivariate “ordering” is established as a pre-ordering $\lesssim$ over $\mathbb{R}^{K}$; a pre-ordering differs from a total ordering in that it need not satisfy the antisymmetric axiom of a total ordering.
+Roughly speaking, an “acceptance region,” so called as it serves as the criterion used to ultimately decide which $y$ are accepted into the prediction region, is then defined as $\widehat{\mathcal{Q}}:=\{s\mid s\lesssim\widehat{q}\}$, where $\widehat{q}$ is the $1-\alpha$ empirical quantile of $\mathcal{S}_{C}$ under $\lesssim$.
+Such a $\widehat{\mathcal{Q}}$ naturally generalizes the standard scalar acceptance interval of $[0,\widehat{q}]$ in the case of non-negative score functions.
+We briefly highlight the distinction between acceptance regions and prediction regions: the former are subsets of the space of multivariate scores, i.e.
+$\subset\mathbb{R}^{K}$, that ultimately define the criteria for retaining particular $y$ values in the prediction region whereas the latter are the subsets of the output space that ultimately have coverage guarantees, i.e.
+$\subset\mathcal{Y}$.
+The two, however, are directly related; in particular, for a fixed score $s(x,y)$, a larger acceptance region will result in a more conservative prediction region.
+Crucially, therefore, the problem of choosing this pre-ordering closely parallels that of choosing $g$, where a poorly chosen pre-ordering will result in overly large acceptance regions and, hence, conservative prediction regions.
+For instance, using a lexicographical ordering $\lesssim_{\mathrm{Lex}}$ will result in axis-aligned hyper-rectangular acceptance regions.
+As a result, rather than manually prescribing a pre-ordering, we define $\lesssim$ in a data-driven fashion by prescribing an indexed family of nested sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, such that $\mathcal{A}_{t_{1}}\subset\mathcal{A}_{t_{2}}$ for $t_{1}\leq t_{2}$ and stating $s_{1}\lesssim s_{2}$ if $\forall t$, $s_{2}\in\mathcal{A}_{t}\implies s_{1}\in\mathcal{A}_{t}$.
+For a family of sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, we take each $\mathcal{A}_{t}$ to be the region of the positive orthant $\mathbb{R}_{+}^{K}$ bounded by the coordinate axes and an “outer frontier” parameterized by $t$.
+The shape of this outer frontier remains fixed over the family and is merely scaled outward from the origin with $t$.
+Under this choice, performing a comparison of $s_{1},s_{2}\in\mathbb{R}^{K}$ in the aforementioned fashion, i.e.
+checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the smallest $t$ for which the outer frontier of $\mathcal{A}_{t}$ intersects $s$.
 Notably, $t(s)$ is precisely the aforementioned data-driven score fusion function $g(s)$ of interest.
-Defining a data-adaptive $g(s)$, therefore, reduces to having a data-driven approach for defining the base contour of $\mathcal{A}_{t}$.
-We additionally restrict this base contour to be a convex set; if such a set were permitted to be nonconvex, computing $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$ would potentially be computationally expensive, increasing the difficulty in defining the resulting prediction regions.
-As demonstrated in Section 3.3, many downstream applications require such efficient computation and, hence, benefit from this restriction to convex sets.
-To have tight quantile regions, we formally wish for the pre-ordering to have the property that the quantile region given by $\widetilde{\mathcal{Q}}:=\{s\mid s\lesssim\widehat{q}\}$, where $\widehat{q}$ is the $1-\alpha$ empirical quantile of $\mathcal{S}_{C}$ under $\lesssim$, should have minimal Lebesgue measure.
-Notably, by construction, such a $\widetilde{\mathcal{Q}}$ will capture $1-\alpha$ points of $\mathcal{S}_{C}$.
-The problem of discovering an optimal pre-ordering can, thus, be equivalently stated as seeking to define the base contour of $\mathcal{A}_{t}$ to match that of the tightest $1-\alpha$ convex cover of $\mathcal{S}_{C}$.
-This final reframing, therefore, naturally motivates selecting the base contour as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$.
-Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangebility of its points with future test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$.
-The goal, then, follows as seeking to define the base contour as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$.
+Defining a data-adaptive $g(s)$, therefore, reduces to having a data-driven approach for defining the outer frontier of $\mathcal{A}_{t}$.
+We restrict this outer frontier to be such that $\mathcal{A}_{t}$ is a convex set; if $\mathcal{A}_{t}$ were permitted to be nonconvex, computing $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$ would potentially be computationally expensive.
+The benefits of such convexity are highlighted, for example, in Section 3.2.
+To have tight acceptance regions, we formally wish for the pre-ordering to have the property that the acceptance region given by $\widetilde{\mathcal{Q}}$ has minimal Lebesgue measure and captures $1-\alpha$ points of $\mathcal{S}_{C}$.
+The problem of discovering an optimal pre-ordering can, thus, be equivalently stated as seeking to define the outer frontier of $\mathcal{A}_{t}$ to match that of the tightest $1-\alpha$ convex cover of $\mathcal{S}_{C}$.
+This final reframing, therefore, naturally motivates selecting the outer frontier to be the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$.
+Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangeability of its points with future test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$.
+The goal, then, follows as seeking to define the outer frontier as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$.
 For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute the final $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$.
 Such a split is predicated on the assumption that the $1-\alpha$ quantile envelope defined over $\mathcal{S}_{C}^{(1)}$ resembles that of $\mathcal{S}_{C}^{(2)}$, implying the $|\mathcal{S}_{C}^{(1)}|$ should be sufficiently large as to capture this structure accurately.
-We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$, starting by selecting the projection directions $\{u_{m}\}$.
+We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$ using a technique paralleling that described in Section 2.2.
+In particular, we start by selecting the projection directions $\{u_{m}\}$ of Equation 1.
 The restriction of $s$ to the positive orthant induces a natural modification of Equation 1, namely where $u\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$ instead of $\mathcal{S}^{K-1}$.
-Exact solution of finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [48].
-If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values of $v$.
+To best approximate Equation 1, we wish for $\{u_{m}\}$ to be uniformly distributed over $\mathcal{S}_{+}^{K-1}$; however, exactly finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [41] If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values of $v$.
 We now wish to define the quantile thresholds $\{\widetilde{q}_{m}\}$ for the selected directions to optimally capture $1-\alpha$ of $\mathcal{S}_{C}^{(1)}$.
 Naively taking the $1-\alpha$ quantile per projection direction $u_{m}$ results in joint coverage by $\widetilde{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$ of $\mathcal{S}^{(1)}_{\mathcal{C}}$ to be $<1-\alpha$.
-To combat this, a straightforward fix is to replace the $1-\alpha$ quantile per direction instead with its Bonferroni-corrected $1-\alpha/M$ quantile.
+A straightforward fix is to replace the $1-\alpha$ quantile per direction instead with its Bonferroni-corrected $1-\alpha/M$ quantile.
 While valid, this approach produces overly conservative prediction regions.
@@
 Assume further that $K$ maps $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined.
-Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\}_{(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}}$.
+Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\mid(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}\}$.
 For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ induced by a collection of nested sets $\{\mathcal{A}_{t}\}_{t\geq 0}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$.
-Then, denoting $\mathcal{C}(X):=\{y:s(X,y)\in\mathcal{Q}(\alpha)\}$, $\mathcal{P}_{X,Y}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha$.
-To compute $\widehat{q}$, for each $s\in\mathcal{S}^{(2)}_{\mathcal{C}}$, we find $t^{*}(s)$, defined to be $\min\{t\in\mathbb{R}:s\in\bigcap_{m=1}^{M}H(u_{m},t\widetilde{q}_{m})\}$.
+Then, denoting $\mathcal{C}(x):=\{y:s(x,y)\in\mathcal{Q}(\alpha)\}$, $\mathcal{P}_{X^{\prime},Y^{\prime}}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha$.
+3.1.2 Score Quantile Threshold
+To then compute $\widehat{q}$, we find $t^{*}(s)$ for each $s\in\mathcal{S}^{(2)}_{\mathcal{C}}$, defined to be $\min\{t\in\mathbb{R}:s\in\bigcap_{m=1}^{M}H(u_{m},t\widetilde{q}_{m})\}$.
 This can be efficiently computed as $t^{*}(s)=\max_{m=1,...,M}(u_{m}^{\top}s/\widetilde{q}_{m})$.
 Denoting the $\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil$-th largest $t^{*}(s)$ as $\widehat{t}$, $\widehat{q}_{m}:=\widehat{t}\widetilde{q}_{m}$ and $\widehat{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widehat{q}_{m})$.
-Note that, if the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$.
-Importantly, while the aforementioned procedure will necessarily result in convex regions $\widehat{\mathcal{Q}}(\alpha)$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, discussed more fully in Section 3.3.
+If the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$.
+Importantly, while the aforementioned procedure will necessarily result in convex regions $\widehat{\mathcal{Q}}$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, as discussed in Section 3.2.
 Such flexibility is necessary, as many studies have demonstrated the need for nonconvex prediction regions for downstream utility, such as [36, 42, 43].
 However, it is unsurprising such flexibility exists, as even a single scalar score $s_{1}(x,y)$ can produce nonconvex prediction regions.
-Another point to emphasis is that this procedure does not require a fusion algorithm to be defined: this is a key conceptual separation between MVCP and traditional conformal prediction.
-In the classic approach, a fused predictor is conformalized; here, instead, the conformalized predictors are fused.
-As a result, there is no notion of a point prediction for MVCP: it only returns prediction regions.
-This is fundamentally unlike the traditional case, where a point prediction $\widehat{y}:=\mathcal{F}(x)$ is “wrapped around” by a surrounding prediction region.
-We present the full algorithm in Algorithm 1 and an accompanying visual walkthrough of the procedure in Appendix A.
-Algorithm 1 MVCP: $\textsc{UnifHypersphere}(K)$ is a subroutine that samples $\sim\mathrm{Unif}(\mathcal{S}^{K-1})$.
+We present the full algorithm in Algorithm 1.
+Algorithm 1 CSA: $\textsc{UnifHypersphere}(K)$ is an assumed subroutine that samples $\sim\mathrm{Unif}(\mathcal{S}^{K-1})$.
 1:  Inputs: Score functions $s_{1},...,s_{K}:\mathcal{X}\rightarrow\mathcal{Y}$, Calibration set $\mathcal{D_{C}}$, Desired coverage $1-\alpha$
-$[\beta_{\mathrm{lo}},\beta_{\mathrm{hi}}]\leftarrow[\alpha/M,\alpha],\widehat{\mathcal{Q}}\leftarrow\emptyset,\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{1}(x_{i},y_{i}),...,s_{K}(x_{i},y_{i}))\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}}$
-3:$\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
-4:while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
-5:     $\beta\leftarrow\frac{\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}}}{2}$
-6:     $\left\{\widetilde{q}_{m}\leftarrow\frac{\lceil(N_{\mathcal{C}_{1}}+1)(1-\beta)\rceil}{N_{\mathcal{C}_{1}}}\text{ quantile of }\{u_{m}^{\top}s_{i}\}_{i=1}^{N_{\mathcal{C}_{1}}}\right\}_{m=1}^{M}$
-7:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
-8:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$
-9:     else $\beta_{\mathrm{hi}}\leftarrow\beta$
-10:$\widehat{t}\leftarrow\frac{\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil}{N_{\mathcal{C}_{2}}}\text{ quantile of }\{\max_{m=1,...,M}(u_{m}^{\top}s_{i}/\widetilde{q}_{m})\}_{i=1}^{N_{\mathcal{C}_{2}}}$
-11:Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
-3.3 MVCP: Prediction Region Definition
+2:  $[\beta_{\mathrm{lo}},\beta_{\mathrm{hi}}]\leftarrow[\alpha/M,\alpha],\widehat{\mathcal{Q}}\leftarrow\emptyset$
+3:  $\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{k}(x_{i},y_{i}))_{k=1}^{K}\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}}$
+4:  $\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
+5:  while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
+6:     $\beta\leftarrow\frac{\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}}}{2}$
+7:     $\left\{\widetilde{q}_{m}\leftarrow(1-\beta)\text{ emp. quantile: }\{u_{m}^{\top}s_{i}\}_{s_{i}\in S_{C}^{(1)}}\right\}_{m=1}^{M}$
+8:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
+9:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$
+10:     else $\beta_{\mathrm{hi}}\leftarrow\beta$
+11:  end while
+12:  $\widehat{t}\leftarrow(1-\alpha)\text{ emp. quantile: }\{\max_{m\in[M]}\frac{u_{m}^{\top}s_{i}}{\widetilde{q}_{m}}\}_{s_{i}\in S_{C}^{(2)}}$
+13:  Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
+3.2 Predict-Then-Optimize
 With this generalization of the score function, a natural question is how to leverage the resulting prediction regions $\mathcal{C}(x)$.
@@
 In the classification setting, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward.
-For a fixed $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the defining hyperplanes after projection.
+For a fixed $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the thresholds $\widehat{q}_{m}$ after projection.
 In the case of regression, however, the prediction region cannot be explicitly constructed in the general case, since $\mathcal{Y}$ contains uncountably many elements.
@@
 Instead, prediction regions in regression settings are often simply used in their implicit form.
-We, therefore, focus on one particular application, namely that of [26] discussed in Section 2.2, and demonstrate the MVCP prediction regions can be leveraged in their framework.
+We, therefore, focus on one particular application, namely that of [36] discussed in Section 2.3, and demonstrate the CSA prediction regions can be leveraged in their framework.
 Such an extension has natural applications to the settings discussed in this original study.
-For instance, the authors demonstrated their method in a robust traffic routing setting with $c$ being predicted traffic from a probabilistic weather model $q(C\mid X)$ for weather covariates $X$.
-A multiview application, therefore, naturally emerges with having separately tailored predictive models, such as another $q_{2}(C\mid X_{2})$ predicting traffic based on historical traffic trends.
-The original study demonstrated that solving this problem in a computationally efficient manner is feasible by leveraging Danskin’s Theorem so long as $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ is efficiently computable for any fixed $w$.
-We focus on demonstrating that this remains the case for MVCP, specifically considering the case where individual view score functions take the form of a generalization of the “GPCP” score from [26].
-Let $q_{k}(C\mid X_{k})$ represent a generative predictor for view $k$ and $\{\widehat{c}_{jk}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X_{k})$ be drawn samples.
+For instance, the authors demonstrated the utility of their method in a robust traffic routing setting with $c$ being predicted traffic from a probabilistic weather model $q(C\mid X)$ for weather covariates $X$.
+An ensembling approach naturally emerges with having multiple predictive models, such as a $q_{2}(C\mid X)$ predicting traffic based instead on historical trends.
+As in Section 3.1, we note that the below described algorithm is better understood with a visual accompaniment, which we provide in Appendix C.
+[36] demonstrated that solving the robust problem variant $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ in a computationally efficient manner is feasible by performing gradient-based optimization on $w$, where the gradient $\nabla_{w}\phi(w)$ of $\phi(w):=\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ can be computed by leveraging Danskin’s Theorem so long as $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ is efficiently computable for any fixed $w$.
+We focus on demonstrating that this remains the case for CSA, specifically considering the case where individual view score functions take the form of the “GPCP” score considered therein.
+In this setup, each constituent predictor is a generative model $q_{k}(C\mid X)$ from which $\{\widehat{c}_{kj}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X)$ samples are drawn.
 Note that $J_{k}$ need not be constant across $k$.
-Then, we define the score to be
-$s_{k}(x,c)=\min_{j}\left[g_{k}\left(\widehat{c}_{jk},c\right)\right],$ (3)
-where $g_{k}(\widehat{c},c)$ is any function that is convex in $c$ for a fixed $\widehat{c}$.
-Notably this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X_{k})=\delta(f_{k}(X_{k}))$.
-To now compute $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, we first let $\vec{j}\in\mathcal{J}=\{j_{1},...,j_{K}\}$ be an indexing tuple, where each $j_{k}\in\{1,...,J_{k}\}$.
-That is, each $\vec{j}$ is a vector that “selects” one sample per view.
+The GPCP score, used to define the score components, is
+$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}\left[d_{k}\left(\widehat{c}_{kj},c\right)\right],$ (3)
+where $d_{k}(\widehat{c},c)$ is a norm over the $C$ space.
+Notably, this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X)=\delta(f_{k}(X))$.
+To compute $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, we first let $\vec{j}\in\mathcal{J}=\{j_{1},...,j_{K}\}$ be an indexing tuple, where each $j_{k}\in\{1,...,J_{k}\}$.
+That is, each $\vec{j}$ is a vector that “selects” one sample per predictor.
 Notably then, the projection $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)$ is convex in $c$, since the projection directions are all restricted to $\mathcal{S}_{+}^{K-1}$.
 As a result, we can compute the maximizer $c^{*}$, since
-$\begin{gathered}c_{\vec{j}}^{*}:=\operatorname*{arg\,max}_{c}f(w,c)\quad\textrm{s.t.}\quad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\forall m\in\{1,...,M\}\end{gathered}$ (4)
+$c_{\vec{j}}^{*}$ $:=\operatorname*{arg\,max}_{c}f(w,c)$ (4)
+   s.t. $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\quad\forall m\in\{1,...,M\}$
 remains a standard convex optimization problem.
 The final maximum can then be found by aggregation, namely $c^{*}=\operatorname*{arg\,max}_{\vec{j}\in\mathcal{J}}f(w,c_{\vec{j}}^{*})$.
-While $|\mathcal{J}|=\prod_{k=1}^{K}J_{k}$, we note that most cases of late-stage fusion in practice tend to have a limited number of views, most typically $K=2$ or $K=3$.
+While $|\mathcal{J}|=\prod_{k=1}^{K}J_{k}$, we note that certain cases of ensemble prediction, such as multi-view prediction, tend to have a limited number of predictors in practice, most typically $K=2$ or $K=3$.
 This coupled with the fact that computing over these indices is trivially parallelizable means this approach is still computationally tractable.
 The full procedure is outlined in Algorithm 2.
-Algorithm 2 Predict-Then-Optimize Under MVCP Prediction Regions
-1:Inputs: Context $x$, CGMs $\{q_{k}(C\mid X_{k})\}_{k=1}^{K}$, Optimization steps $T$, Sample counts $\{J_{k}\}_{k=1}^{K}$, MVCP quantile $\widehat{\mathcal{Q}}:=\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$
-2:$w\sim U(\mathcal{W}),\{\{\widehat{c}_{jk}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X_{k})\}_{k=1}^{K},\mathcal{J}=\prod_{k=1}^{K}[J_{k}]$
-3:for $t\in\{1,\ldots T\}$ do
-4:     $\left\{c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}\left\{f(w,c)\mid u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\forall m\in\{1,...,M\}\right\}\right\}_{\vec{j}\in\mathcal{J}}$
-5:     $c^{*}\leftarrow\operatorname*{arg\,max}_{c_{\vec{j}}^{*}}f(w,c_{\vec{j}}^{*})$
-6:     $w\leftarrow\Pi_{\mathcal{W}}(w-\eta\nabla_{w}f(w,c^{*}))$
-7:Return $w$
+Algorithm 2 Predict-Then-Optimize Under CSA
+1:  Inputs: Context $x$, Predictors $\{q_{k}(C\mid X)\}_{k=1}^{K}$, Optimization steps $T$, Sample counts $\{J_{k}\}_{k=1}^{K}$, CSA quantile $\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$
+2:  $\{\{\widehat{c}_{kj}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X)\}_{k=1}^{K},\mathcal{J}=\prod_{k=1}^{K}[J_{k}]$
+3:  $w^{(0)}\sim U(\mathcal{W})$
+4:  for $t\in\{1,\ldots T\}$ do
+5:     for $\vec{j}\in\mathcal{J}$ do $c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}f(w^{(t)},c)$$\mathrm{s.t.}\quad\forall m\in 1,...,M\qquad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}$
+6:     $c^{*}\leftarrow\operatorname*{arg\,max}_{c_{\vec{j}}^{*}}f(w^{(t)},c_{\vec{j}}^{*})$
+7:     $w^{(t)}\leftarrow\Pi_{\mathcal{W}}(w^{(t-1)}-\eta\nabla_{w}f(w^{(t-1)},c^{*}))$
+8:  end for
+9:  Return $w^{(T)}$
+Table 1: Average coverages across tasks for $\alpha=0.10$ are shown in the top rows and average prediction set sizes in the bottom rows.
+Both were assessed over a batch of i.i.d.
+test samples (15% of the validation set from ImageNet).
+Standard deviations and means were computed across 10 randomized draws of the calibration and test sets.
+Dataset Models ResNet VGG DenseNet VFCP $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA
+ImageNet Coverage 0.901 (0.005) 0.902 (0.003) 0.902 (0.003) 0.899 (0.004) 0.938 (0.003) 0.909 (0.004) 0.9 (0.004) 0.9 (0.003)
+Size 137.004 (1.98) 136.116 (2.206) 120.096 (2.427) 46.063 (1.089) 87.337 (1.604) 82.746 (1.692) 131.856 (2.378) 34.006 (0.924)
 4 Experiments
-We now study MVCP empirically across several tasks, demonstrating its improvement over alternate approaches in a synthetic classification task (Section 4.1), in predict-then-optimize tasks (Section 4.2), and in a classification task over a non-synthetic dataset of interest, namely Handwritten [50] (Section 4.3).
+We now study CSA empirically across several tasks, demonstrating its coverage guarantees with reduced conservatism.
+We demonstrate improvements in an ImageNet classification task in Section 4.1, in real-data UCI regression tasks as studied in [29] in Section 4.2, and in a downstream predict-then-optimize task in Section 4.3.
 To allow for fair comparison, we note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment.
-This in turn means that some care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for MVCP, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction.
+This in turn means that some care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for CSA, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction.
 For this reason, we fixed the split to be $20\%$-$80\%$ in experiments.
-Further details are provided in Appendix E, and code is available at: https://github.com/yashpatel5400/fusioncp
-4.1 Synthetic Classification Task
-We first consider a synthetic classification task, fitting multinomial logistic regression models $\widehat{f}_{k}:\mathcal{X}_{k}\rightarrow\mathcal{Y}$, where $\mathcal{X}=\mathbb{R}^{120}$ and $\mathcal{Y}=\{1,...,20\}$.
-We consider $K$-view classification problems, with $K\in\{2,...,6\}$, where the views here partition the input dimension into subspaces of dimension $120/K$, namely $\mathcal{X}_{i}=\mathbb{R}^{1+(d/K)(i-1):(d/K)i}$ for $i\in\{1,...,K\}$.
-We compare MVCP to four baselines: a fusion-free approach that conformalizes the view with the highest predictive accuracy; an approach using a predefined “Norm” fusion algorithm as described in Section 3.1, namely with $g_{1}(s)=(\sum_{k=1}^{K}s_{k}^{2})^{1/2}$; another similarly using a predefined “Sum” fusion algorithm with $g_{2}(s)=\sum_{k=1}^{K}s_{k}$; and one that conformalizes a fused predictor as also discussed in Section 3.1, with $\mathcal{F}=(\widehat{f}_{1}(x)+\dots+\widehat{f}_{K}(x))/K$.
-We perform conformalization using the standard classification score function across all approaches, namely $s_{k}(x,y)=\sum_{j=1}^{l}\widehat{f}_{k}(x)_{\pi_{j}(x,k)}$ where $y=\pi_{l}(x,k)$ and $\pi(x,k)$ is the permutation of $\{1,\ldots,20\}$ that sorts $\widehat{f}_{k}\left(x\right)$ from most likely to least likely.
+We herein compare against the methods presented in Section 2.4, namely the model selection of [27], the aggregation methods of [25], and the single weighted score projection (VFCP) of [40].
+We specifically consider the following aggregation methods from [25]: the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches, fully described in Appendix F.
+Notably, these methods do not lend themselves for use in the predict-then-optimize setting, so we eliminate them from consideration therein.
+Similarly, VFCP can only be applied in classification settings; we, thus, do not compare to it across the regression tasks.
+Code will be made public upon acceptance.
+4.1 Classification Tasks
+We first study the predictive efficiency of the aforementioned methods on the ImageNet classification task [44].
+In particular, an ensemble was constructed from three separately trained deep learning architectures, namely ResNet-50, VGG-11, and DenseNet-121.
+Conformalization on the individual models was performed using the standard classification score function across all approaches, namely $s(x,y)=\sum_{j=1}^{l}\widehat{f}(x)_{\pi_{j}(x)}$ where $y=\pi_{l}(x)$ and $\pi(x)$ is the permutation of $\{1,\ldots,|\mathcal{Y}|\}$ that sorts $\widehat{f}\left(x\right)$ from most to least likely.
+Calibration was performed using 85% of the ImageNet test set and assessment of the coverage and interval lengths on the remaining 15%, with 10 trials conducted over randomized draws of these calibration and test sets.
 The results are presented in Table 1.
-Table 1: Coverages across tasks for $\alpha=0.10$ are shown in the left table, where coverage was assessed over a batch of 300 i.i.d.
+We see that all the approaches exhibit the desired coverage guarantees.
+However, the CSA score approach consistently produce significantly smaller prediction regions than both the individually conformalized input views and alternate aggregation strategies.
+4.2 UCI Regression Tasks
+We now similarly study the predictive efficiency of CSA across a suite of regression tasks from the UCI repository [45] as studied by [29].
+We specifically consider those tasks for which there are at least 1,000 samples; the complete list of tasks are provided in Appendix D.
+The data for each task were split with 80% for training, 10% for calibration, and 10% for testing coverage and interval lengths.
+The problem setup was replicated from [29], in which four prediction methods were ensembled for each prediction task, namely an OLS model, a LASSO linear model, a random forest, and a multi-layer perceptron.
+A standard absolute value residual function was used as the score across all methods, namely $s(x,y)=|\widehat{f}(x)-y|$.
+Prediction intervals could be explicitly constructed for the $\mathcal{C}^{M}$, $\mathcal{C}^{R}$, and $\mathcal{C}^{U}$ aggregation methods.
+For assessing interval length for CSA, however, a discretized grid $\mathcal{G}_{Y}\subset\mathcal{Y}$ of coarseness $\Delta y$ was considered, and an interval length estimate given by $\mathcal{L}(\mathcal{C}(x))\approx\Delta y\cdot|\{y:y\in\mathcal{G}_{Y},s(x,y)\in\widehat{\mathcal{Q}}\}|$.
+$10$ trials were conducted for each task, randomized over subset selections of the training, calibration, and test sets.
+We also present an ablation, labeled “CSA (Single-Stage),” to demonstrate the two-stage calibration described in Section 3.1 is necessary to retain coverage; this single-stage approach does not split $\mathcal{S}_{C}$ and instead directly computes $\{\widehat{q}_{m}\}$ on $\mathcal{S}_{C}$ per Section 3.1.1.
+A subset of the results are given in Table 2; the full set of results is deferred to Appendix D due to space limitations.
+As in the results presented in Section 4.1, we see that CSA retains the coverage guarantees typical of conformal prediction yet produces significantly smaller prediction intervals than both the individual models and the alternate aggregation strategies.
+We additionally see that the “CSA (Single-Stage)” approach fails to retain coverage, demonstrating the necessity of the two-stage calibration.
+Table 2: Average coverages across tasks for $\alpha=0.05$ are shown in the top rows and average prediction set lengths in the bottom rows, both were assessed over a batch of i.i.d.
+test samples (10% of the datasets).
+Standard deviations and means were computed across 10 randomized draws of the training, calibration, and test sets.
+Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, thus, precluded from comparison.
+Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
+3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
+Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
+parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
+Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
+pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
+Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
+pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
+Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
+slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
+Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
+4.3 CSA Predict-Then-Optimize
+We now study the use of CSA for a real-world predict-then-optimize traffic routing task, as studied in the conformal predict-then-optimize paper [36].
+In this task, a time series of $T$ previously observed precipitation readings are used to predict future precipitations and, in turn, anticipated traffic, as fully described in Appendix E.
+We consider the traffic routing problem for a fixed source-target pair $(s,t)$ over the graph of Manhattan, where $|\mathcal{V}|=4584$ and $|\mathcal{E}|=9867$.
+Formally, such a problem seeks
+$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}\widehat{c}^{T}w$ (5)
+[…]
+where $x\in\mathbb{R}^{T\times H\times W}$ are the previous precipitation readings, $w_{e}\in\mathbb{R}^{|\mathcal{E}|}$ the traffic proportion routed along road $e$, $c\in\mathbb{R}^{|\mathcal{E}|}$ the transit times anticipated across roads, $A\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{E}|}$ the graph incidence matrix, and $b\in\mathbb{R}^{|\mathcal{V}|}$ the vector that specifies the routing problem, in which $b_{s}=1,b_{t}=-1,$ and $b_{k}=0$ for $s$ the travel source node, $t$ the terminal node, and $k\notin\{s,t\}$ all other nodes.
+We then consider two probabilistic models for traffic prediction, namely one based on the classical probabilistic Lagrangian integro-difference approach (STEPS) of [46] and one on the modern latent diffusion model (LDM) approach of [47].
+As a result of the higher inference cost of the latter, we consider the setup where $J_{1}>J_{2}$, specifically with $J_{1}=4$ and $J_{2}=1$, which additionally serves to empirically highlight the flexibility of non-uniform sampling from predictors discussed in Section 3.2.
+As discussed in Section 4, the alternate aggregation strategies do not lend themselves for use in this setting.
+We, therefore, only compare CSA to the best model selection strategy, specifically by comparing to the separate conformalizations of the two constituent predictors.
+Each model was conformalized with Equation 3 taken to be the score, where $d_{k}$ was set to be the standard 2-norm across both predictors, such that $s_{1}(x,c)=\min_{j\in 1,...,J_{1}}||\widehat{c}_{1j}-c||_{2}$ for $q_{1}(C\mid x)$ and similarly for $q_{2}(C\mid x)$.
+We here evaluate the methods using the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.3.
+This quantity measures the conservatism of the produced robust optimal value and is bounded in $[0,1]$ in this setting.
+Experiments were conducted with $|\mathcal{D}_{C}|=200$.
+The suboptimality was then computed across 100 i.i.d.
 test samples.
-Average prediction set sizes are shown in the right table, averaged over a batch of 300 i.i.d.
-test samples with standard deviations in parentheses.
-$K$ Best View Pre-Fused Sum Norm MVCP
-2 0.89 0.91 0.91 0.91 0.90
-3 0.89 0.91 0.91 0.91 0.90
-4 0.89 0.91 0.90 0.90 0.91
-5 0.88 0.90 0.90 0.90 0.90
-6 0.89 0.91 0.90 0.90 0.90
-Best View Pre-Fused Sum Norm MVCP
-10.84 (0.34) 8.11 (0.56) 8.73 (0.42) 8.72 (0.42) 7.58 (0.42)
-12.36 (0.30) 8.49 (0.44) 9.05 (0.32) 8.83 (0.35) 7.68 (0.42)
-13.32 (0.30) 8.75 (0.51) 9.04 (0.47) 8.41 (0.50) 8.07 (0.50)
-13.90 (0.19) 9.01 (0.59) 8.91 (0.63) 8.45 (0.59) 8.37 (0.58)
-14.28 (0.27) 9.26 (0.39) 9.04 (0.68) 8.71 (0.71) 8.71 (0.75)
-We see that all the approaches exhibit the desired coverage guarantees.
-However, the multivariate score approaches consistently produce significantly smaller prediction regions than the individually conformalized input views, with MVCP consistently producing the smallest sets on average.
-4.2 MVCP Predict-Then-Optimize
-We now study the use of MVCP across predict-then-optimize tasks, specifically considering the fractional knapsack problem as in the original conformal predict-then-optimize paper [26]:
-$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(X)}\quad-\widehat{c}^{T}w$ (5)
-[…]
-where $p\in\mathbb{R}^{n}$ and $B>0$.
-The distributions $\mathcal{P}(C)$ and $\mathcal{P}(X\mid C)$ are again taken to be those from various simulation-based inference (SBI) benchmark tasks provided by [51], chosen as they have $\mathcal{P}(C\mid X)$ with complex structure, fully described in Appendix C.
-We specifically study the Gaussian Mixture distributional setup across increasing dimensions to highlight the capacity for MVCP to succeed even for complex combinations of distributions.
-We then consider $K=2$ views, namely with $\mathcal{X}=\mathbb{R}^{d}$, $\mathcal{X}_{1}=\mathbb{R}^{1:d_{1}}$, and $\mathcal{X}_{2}=\mathbb{R}^{d_{1}+1:d}$.
-For each $\mathcal{X}_{k}$, we train a generative model $q_{k}(C\mid X_{k})$, specifically a normalizing flow.
-We then compare four distinct ways of formulating $\mathcal{C}(X)$, namely by conformalizing each view individually, by conformalizing a fused map, by fusing the scores via a sum, and by fusing with MVCP, all using the score given by Equation 3, specifically taking $g_{k}$ to be the $\ell^{2}$ metric.
-The “Pre-fused” strategy employed here is again a mean aggregation, i.e.
-a sample $\widehat{c}\sim\mathcal{F}(q_{1},q_{2})$ is drawn via $\widehat{c}_{1}\sim q_{1}(C\mid X_{1}),\widehat{c}_{2}\sim q_{2}(C\mid X_{2})$ and $\widehat{c}$ defined as $(\widehat{c}_{1}+\widehat{c}_{2})/2$.
-As discussed in Section 3.3, the prediction regions are not explicitly constructed in regression settings, rendering the reporting of prediction region sizes impossible.
-We instead compare the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.2.
-This quantity measures the conservatism of the produced robust optimal value and is bounded in $[0,1]$ in this setting, with the upper bound seen as Equation 5 can trivially be solved by simply taking $w=0$.
-The results are presented in Table 2.
-Table 2: Coverages across tasks for $\alpha=0.05$ are shown in the left table, where coverage was assessed over a batch of 500 i.i.d.
+To, therefore, assess the improvement offered by CSA, we conduct two paired t-tests, with hypotheses
+$H_{0}:\Delta^{(\mathrm{CSA})}_{\%}=\Delta^{(\mathrm{STEPS})}_{\%}\quad H_{1}:\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$
+and similarly for $\Delta^{(\mathrm{CSA})}_{\%}$ and $\Delta^{(\mathrm{LDM})}_{\%}$.
+The coverages and results of such t-tests are provided in Table 3, from which we find that CSA significantly reduces the suboptimality after accounting for Bonferroni multiple testing.
+We see that, while conformalization of either of the two views individually already produces the desired coverage, CSA produces more informative prediction regions, and hence less conservative robust upper bounds.
+Table 3: Coverages across tasks for $\alpha=0.05$ are shown in the left table across both the individually conformalized and CSA fused approaches.
+P-values of the paired t-tests comparing gap proportions ($\Delta_{\%}$) are shown in the right table.
+Both were computed over 100 i.i.d.
 test samples.
-Suboptimality gap proportions ($\Delta_{\%}$) are shown in the right table, with medians reported over a batch of 500 i.i.d.
-test samples and median average deviations in parentheses.
-View 1 View 2 Pre-Fused Sum MVCP
-Gaussian Mixture (1) 0.964 0.958 0.952 0.964 0.976
-Gaussian Mixture (2) 0.938 0.938 0.93 0.95 0.94
-Gaussian Mixture (3) 0.944 0.972 0.936 0.944 0.948
-Gaussian Mixture (4) 0.942 0.94 0.914 0.95 0.924
-View 1 View 2 Pre-Fused Sum MVCP
-0.901 (0.1003) 0.886 (0.1144) 1.0 (0.0) 0.702 (0.1403) 0.542 (0.0946)
-0.962 (0.0453) 1.0 (0.0155) 1.005 (0.0019) 0.893 (0.1084) 0.874 (0.0792)
-1.011 (0.018) 0.997 (0.0208) 1.011 (0.0035) 0.838 (0.0735) 0.868 (0.047)
-1.026 (0.0181) 0.987 (0.0327) 1.018 (0.0066) 1.003 (0.0144) 0.946 (0.0286)
-We see that all methods produce the coverage guarantees in accordance to Theorem 3.1.
-The resulting suboptimality gaps additionally exhibit two notable characteristics.
-The first is that MVCP produces more informative prediction regions, and hence less conservative robust upper bounds, if the views in consideration are individually informative, that is, each has $\Delta_{\%}<1$.
-While the sum conformal fusion exhibits a similar pattern, its effective reduction is consistently less than that provided by MVCP.
-Additionally, in cases where one of the two views is uninformative, i.e.
-$\Delta_{\%}=1$, MVCP nearly retains the performance of the informative view, as the quantile envelope adapts in considering all projection directions.
-The sum fusion technique, on the other hand, significantly degrades from the performance of the optimal view.
-4.3 Real Data Classification
-We now conduct experiments on the Handwritten dataset [50].
-This dataset consists of features of 2000 handwritten digits $(0-9)$ (200 examples per class) with 6 different views of each image extracted from a collection of Dutch utility maps, described fully in Appendix D.
-The baseline approaches parallel those presented in Section 4.1.
-The results are compiled in Table 3.
-As in the synthetic classification experiment, we see that considering multiple views produces significantly more informative prediction regions than solely conformalizing a single view, especially when done in the manner prescribed by MVCP.
-Table 3: Coverages across tasks for $\alpha=0.10$ are shown in the left table, where coverage was assessed over a batch of 295 i.i.d.
-test samples.
-Average prediction set sizes are shown in the right table, averaged over a batch of 295 i.i.d.
-test samples with standard deviations in parentheses.
-$K$ Best View Pre-Fused Sum Norm MVCP
-2 0.90 0.91 0.91 0.91 0.90
-3 0.90 0.90 0.90 0.90 0.90
-4 0.91 0.90 0.90 0.90 0.90
-5 0.91 0.91 0.91 0.90 0.90
-6 0.90 0.91 0.90 0.90 0.91
-Best View Pre-Fused Sum Norm MVCP
-3.39 (0.09) 3.69 (0.11) 3.52 (0.13) 3.51 (0.13) 2.68 (0.17)
-3.39 (0.09) 3.48 (0.09) 2.97 (0.09) 2.95 (0.09) 1.84 (0.10)
-3.28 (0.14) 3.42 (0.05) 2.87 (0.05) 2.83 (0.05) 1.64 (0.12)
-3.28 (0.14) 3.12 (0.07) 2.73 (0.08) 2.68 (0.07) 1.55 (0.06)
-3.19 (0.17) 3.07 (0.11) 2.71 (0.06) 2.64 (0.05) 1.60 (0.07)
+STEPS LDM CSA
+0.981 0.962 0.968
+$H_{1}$ p-value
+$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$ $3.61\times 10^{-4}$
+$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{LDM})}_{\%}$ $9.50\times 10^{-4}$
 5 Discussion
-We have presented MVCP, a framework for producing informative prediction regions in late-stage fusion pipelines.
+We have presented CSA, a framework for producing informative prediction regions in ensemble predictor pipelines.
 We additionally demonstrated the generality of the proposed pipeline in settings of both classification and robust predict-then-optimize regression tasks.
 This work suggests many directions for extension.
-One point of interest is the integration of MVCP with other downstream regression applications.
-For instance, [52] proposed an end-to-end differentiable extension to [26]: extending MVCP for integration to their framework would be of great interest.
-Additionally, given the prevalence of late-stage fusion in robotics applications, an exciting applied avenue of research would investigate the optimization of MVCP to leverage it in settings of robust control.
-6 Acknowledgements
-We acknowledge the support of the College of Literature, Science, and the Arts at the University of Michigan via the Meet the Moment Research Initiative.
-Appendix A Multivariate Visualization
+One point of interest is the integration of CSA with other downstream regression applications.
+For instance, [48] proposed an end-to-end differentiable extension to [36]: extending CSA for integration to their framework would be of great interest.
+Additionally, given the prevalence of late-stage fusion in robotics applications, an exciting applied avenue of research would investigate the use of CSA in settings of robust control.
+Appendix A CSA Visual Walkthrough
 We walk through a visual presentation of the approach below to supplement the textual description in the main text.
-As discussed, we start with a collection of multivariate calibration scores $\mathcal{S}_{C}$ with $s\in\mathcal{S}$ being $\in\mathbb{R}^{K}$.
+We start with a collection of multivariate calibration scores $\mathcal{S}_{C}$ with $s\in\mathcal{S}$ being $\in\mathbb{R}^{K}$.
 For the purposes of visualization in this section, we have $K=2$.
-We first partition the score evaluations $\mathcal{S_{C}}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$ into a subset used to define the pre-ordering and the remainder to define the multivariate quantile.
-Figure 2: The calibration score evaluations are first split between those used to define the pre-ordering (green) and those used to define the final multivariate quantile (red).
+We first partition the score evaluations $\mathcal{S_{C}}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, with a subset $\mathcal{S}_{C}^{(1)}$ used to define the pre-ordering and the remainder $\mathcal{S}_{C}^{(2)}$ to define the multivariate quantile.
+Figure 2: The calibration score evaluations are first split between those used to define the pre-ordering (green) $\mathcal{S}_{C}^{(1)}$ and those used to define the final multivariate quantile (red) $\mathcal{S}_{C}^{(2)}$.
 We first wish to define the pre-ordering over $\mathcal{S}_{C}^{(1)}$.
-As described in the main text, the goal is to define this using an indexed family of sets $\mathcal{A}_{t}$ with index $t\in\mathbb{R}$, after which the multivariate quantile reduces to the univariate quantile formulation.
+As described in the main text, the goal is to define this using an indexed family of sets $\mathcal{A}_{t}$ with index $t\in\mathbb{R}$, after which the multivariate quantile approach reduces to the univariate quantile formulation.
 To ensure the final envelope over $\mathcal{S}_{C}^{(2)}$ remains as tight as possible, we wish to define this family in a data-driven fashion.
@@
 Suppose $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$ are exchangeable.
-Assume further that $K$ maps $s_{j}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined.
-Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\}_{(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}}$.
-For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$.
+Assume further that $K$ maps $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined.
+Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\mid(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}\}$.
+For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ induced by a collection of nested sets $\{\mathcal{A}_{t}\}_{t\geq 0}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$.
 Then, denoting $\mathcal{C}(x):=\{y:s(x,y)\in\mathcal{Q}(\alpha)\}$,
-$\mathcal{P}_{X,Y}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha.$ (6)
+$\mathcal{P}_{X^{\prime},Y^{\prime}}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha$ (6)
 Proof.
@@
 completing the proof as desired. ∎
-Appendix C Simulation-Based Inference Benchmarks
-The benchmark tasks are a subset of those provided by [53].
-For convenience, we provide brief descriptions of the tasks curated by this library; however, a more comprehensive description of these tasks can be found in their manuscript.
-C.1 Gaussian Mixture
-A mixture of two Gaussians, with one having a much broader covariance structure:
-$\text{{Prior}: }\beta\sim\mathcal{U}(-10,10)$
-$\text{{Simulator}: }x\mid w\sim 0.5\mathcal{N}(x\mid w,I)+0.5\mathcal{N}(x\mid w,.01I)$
-For any Gaussian Mixture ($k$) distribution referenced in Section 4, we combined separate instances of this setup, where the typical setup has $\beta\in\mathbb{R}^{2}$.
-In other words, we generate $x:=[x_{1},...,x_{k}]$ by drawing a prior over a similarly partitioned $\theta:=[\theta_{1},...,\theta_{k}]$ and taking $x_{i}\sim\mathcal{P}(X_{i}\mid\theta_{i})$ according to the above distribution.
-Appendix D Real Data Featurization
-The data features of the Handwritten dataset [50] are as follow:
-  • 47 Zernike moments
-  • 76 Fourier coefficients of the character shapes
-  • 6 morphological features
-  • 240 pixel averages in $2\times 3$ windows
-  • 64 Karhunen-Love coefficients
-  • 216 profile correlations
-Appendix E Training Details
-All encoders were implemented in PyTorch [54] with a Neural Spline Flow architecture.
-The NSF was built using code from [55].
-Specific architecture hyperparameter choices were taken to be the defaults from [55] and are available in the code.
-Optimization was done using Adam [56] with a learning rate of $10^{-3}$ over 5,000 training steps.
-Minibatches were drawn from the corresponding prior $\mathcal{P}(Y)$ and simulator $\mathcal{P}(X\mid Y)$ as specified per task in the preceding section.
-Training these models required between 10 minutes and two hours using an Nvidia RTX 2080 Ti GPUs for each of the SBI tasks.
+Appendix C CSA Predict-Then-Optimize Visual Walkthrough
+We now present a visual accompaniment of the predict-then-optimize algorithm presented in Section 3.2.
+We once again take $K=2$ for visual clarity in this walkthrough, where the predictors are as discussed in Section 3.2, namely assumed to be generative predictors $q_{k}(C\mid X)$ where the number of samples per predictor are fixed to be $\{J_{k}\}$.
+For illustration, we assume $J_{1}=5$ and $J_{2}=3$, meaning predictions with the first model are made by drawing 5 samples and 3 for the second.
+We assume the CSA calibration of Section 3.1 has already been performed, from which a collection of projection directions and quantiles $\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$ are available that implicitly define an acceptance region $\widehat{\mathcal{Q}}$.
+We further assume the individual predictor score functions are all the GPCP score given in Equation 3, with $d_{k}$ from Equation 3 specifically here taken to simply be the standard Euclidean 2-norm, giving
+$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}||\widehat{c}_{kj}-c||.$ (10)
+We now wish to compute $c^{*}=\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$.
+To do so, we must start by defining this region $\mathcal{C}(x)$ for the test point $x$, which we do by drawing the respective number of samples from the two models, producing samples $\{\widehat{c}_{1j}\}_{j=1}^{5}$ and $\{\widehat{c}_{2j}\}_{j=1}^{3}$, as shown in Figure 7.
+Figure 7: Samples drawn from the two generative models $\{\widehat{c}_{1j}\}_{j=1}^{5}\sim q_{1}(C\mid x)$ (blue) and $\{\widehat{c}_{2j}\}_{j=1}^{3}\sim q_{2}(C\mid x)$ (green).
+Note that this is a visualization in the $\mathcal{C}$ space, i.e.
+not the space of multivariate scores.
+By definition, $\forall c\in\mathcal{C}(x)$,
+$u_{m}^{\top}\left(\min_{j_{1}=1,...,5}||\widehat{c}_{1j_{1}}-c||,\min_{j_{2}=1,...,3}||\widehat{c}_{2j_{2}}-c||\right)\leq\widehat{q}_{m}\qquad\forall m=1,...,M.$ (11)
+As a result, we must have that, $\forall c\in\mathcal{C}(x)$, $\exists$ $j_{1}=1,...,5$ and $j_{2}=1,...,3$ such that $u_{m}^{\top}\left(||\widehat{c}_{1j_{1}}-c||,||\widehat{c}_{2j_{2}}-c||\right)\leq\widehat{q}_{m}$ $\forall m=1,...,M$.
+Solving for $c^{*}$, therefore, amounts to considering each pair $\vec{j}:=(j_{1},j_{2})\in\mathcal{J}$, where $\mathcal{J}:=\{1,...,5\}\times\{1,...,3\}$, and solving
+$c_{\vec{j}}^{*}$ $:=\operatorname*{arg\,max}_{c}f(w,c)$ (12)
+   s.t. $u_{m}^{\top}\left(||\widehat{c}_{1\vec{j}_{1}}-c||,||\widehat{c}_{2\vec{j}_{2}}-c||\right)\leq\widehat{q}_{m}\quad\forall m\in\{1,...,M\}$
+Notice that, for any fixed $\vec{j}$, this is a standard convex optimization problem with a convex feasible region.
+We illustrate how the feasible region would be constructed for a fixed $\vec{j}$ in Figure 8.
+Note that the construction of this feasible is never explicitly done in practice and is only implicitly used by convex solver routines in practice.
+Figure 8: A candidate $c\in\mathcal{C}$ is in prediction region if the projections of its distances $(d_{1},d_{2})$ from at least one pair of points indexed by $\vec{j}:=(j_{1},j_{2})$ is in the acceptance region $\widehat{\mathcal{Q}}$.
+Here, we illustrate a point $c$ that lies in the feasible region from its proximity to the $\vec{j}:=(4,1)$ pair of points.
+Note that the left is again a visualization over the $\mathcal{C}$ space, whereas the right is of the score space.
+We can, therefore, then solve Equation 12 over all possible $\vec{j}\in\mathcal{J}$ and aggregate the maxima to compute $c^{*}$.
+Appendix D Additional UCI Results
+We consider those regression tasks from the UCI repository [45] that have at least 1,000 samples. The complete list of tasks considered herein is provided in Table 4.
+Table 4: Tasks considered for empirical validation of prediction region efficiency and coverage, along with their relevant dimensions. All tasks are scalar regression tasks, i.e. $y\in\mathbb{R}$.
+Dataset Observations Input Dimension
+3droad 434874 3
+airfoil 1503 5
+bike 17379 17
+parkinsons 5875 20
+pol 15000 26
+pumadyn32nm 8192 32
+slice 53500 385
+The complete collection of results is presented in Table 5.
+As in the main text, we see reduced conservatism of the CSA prediction regions compared to alternatives.
+We additionally notice that the single-stage CSA approach fails to retain the coverage guarantees.
+Table 5: Average coverages across tasks for $\alpha=0.05$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d.
+test samples (10% of the dataset size).
+Standard deviations and means were computed across 10 randomizations of draws of the training, calibration, and test sets.
+Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, therefore, precluded from comparison.
+Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
+3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
+Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
+airfoil Coverage 0.95 (0.012) 0.981 (0.018) 0.954 (0.02) 0.969 (0.014) 0.962 (0.014) 0.939 (0.024) 0.969 (0.014) 0.854 (0.04) 0.952 (0.006)
+Length 18.742 (0.313) 25.218 (1.05) 7.353 (0.672) 25.256 (1.424) 18.097 (0.565) 11.96 (0.634) 18.303 (0.787) 1.873 (0.21) 2.714 (0.141)
+bike Coverage 0.963 (0.004) 0.959 (0.004) 0.948 (0.001) 0.955 (0.003) 0.955 (0.003) 0.932 (0.008) 0.959 (0.003) 0.946 (0.005) 0.955 (0.003)
+Length 3.397 (0.131) 3.416 (0.123) 0.025 (0.0) 1.716 (0.104) 1.654 (0.11) 0.837 (0.055) 2.094 (0.09) 0.569 (0.097) 0.629 (0.105)
+parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
+Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
+pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
+Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
+pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
+Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
+slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
+Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
+Appendix E Robust Traffic Routing Setup
+We replicate the experimental setup of [36], namely where a graph of Manhattan with corresponding nominal transit times was extracted using OSMnx [49].
+Formally, the Manhattan graph is given as a tuple $(\mathcal{V},\mathcal{E})$, where the edge weights represent the transit times along the respective city roads.
+Such weights were assigned in a two-step process, namely by first making weather predictions and then using such weather predictions to then upweight the nominal transit times.
+In particular, precipitation forecasts were made from time-series observations of previous precipitations readings, specifically given over a map spatially resolved to $H\times W$ resolution.
+Precipitation forecasters, such as those considered in the experiments herein as given in [46] and [47], specifically map such previous observations to potential future trajectories.
+Formally, they define probabilistic models over some future time horizon $T_{f}$, from which probabilistic draws $\widetilde{Y}\in\mathbb{R}^{T_{f}\times H\times W}\sim\mathcal{P}(\widetilde{Y}\mid x)$ can be made, where $x\in\mathbb{R}^{T\times H\times W}$.
+Notably, we instead consider the probabilistic forecasts at some future fixed time point $T^{\prime}$, meaning the outcome of interest $Y\in\mathbb{R}^{H\times W}=\widetilde{Y}_{T^{\prime}}$
+From a precipitation map, namely a spatially resolved reading $Y\in\mathbb{R}^{H\times W}$, we assign the final edge weights by first associating nodes to the closest pixel coordinate of the precipitation map.
+That is, denoting the pixel nearest to a vertex $v$ as $(p^{v}_{x},p^{v}_{y})$, the node is assigned the value at such a spatial location $Y_{p^{v}_{y},p^{v}_{x}}$.
+To, therefore, assign the edge weight, we average the weights of the edge endpoints and then weigh the nominal transit time.
+In particular, denoting the nominal transit time along such an edge $e$ between nodes $(s,t)$ as $\widetilde{c}_{e}$, the transit time with traffic was computed as
+$c_{e}:=\widetilde{c}_{e}\cdot\exp\left\{\frac{Y_{p^{e_{s}}_{x},p^{v}_{y}}+Y_{p^{e_{t}}_{x},p^{e_{t}}_{y}}}{2}\right\}.$ (13)
+Appendix F Conformal Aggregation Methods
+We now describe the methods from [25] that were compared against experimentally, specifically the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches.
+As discussed in Section 2.4, these methods all follow the structural form of Equation 2 and largely differ in their choice of weights and thresholds.
+The standard majority-vote $\mathcal{C}^{M}$ is the most natural choice, defined by
+$\mathcal{C}^{M}(x):=\left\{y\mid\frac{1}{K}\sum_{k=1}^{K}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]>\frac{1}{2}\right\}.$ (14)
+The randomized methods differ in that independent randomization is leveraged over the threshold, namely with:
+$\mathcal{C}^{R}(x):=\left\{y\mid\frac{1}{K}\sum_{k=1}^{K}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]>\frac{1}{2}+\frac{U}{2}\right\}$ (15)
+$\mathcal{C}^{U}(x):=\left\{y\mid\frac{1}{K}\sum_{k=1}^{K}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]>U\right\},$ (16)
+for $U\sim\mathrm{Unif}([0,1])$. Notably, all these methods retain the guarantees typical of conformal prediction.

### v3

@@
 Methods for conformal aggregation have in turn been proposed for ensembled prediction, where the prediction regions of individual models are merged as to retain coverage guarantees while minimizing conservatism.
-Merging the prediction regions directly, however, sacrifices structures present in the conformal scores that can further reduce conservatism.
+Merging the prediction regions directly, however, can miss out on opportunities to further reduce conservatism by exploiting structures present in the conformal scores.
 We, therefore, propose a novel framework that extends the standard scalar formulation of a score function to a multivariate score that produces more efficient prediction regions.
@@
 By using conformal prediction on a user-defined score function, prediction regions attain marginal coverage guarantees.
-While calibration is guaranteed from this procedure, predictive efficiency, quantified as the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
+While calibration is guaranteed from this procedure, predictive efficiency, i.e.
+the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
 As a result, methods have arisen to perform conformal model aggregation, which both provide uncertainty estimates of the ensembled predictions and do so in ways as to minimize the prediction region size [25, 26, 27, 28, 29].
@@
 Conformal prediction is one such method [23, 24].
-Due to its lack of structural assumptions, conformal prediction pairs well with the increasingly common black-box style of modeling.
 Conformal prediction serves as a wrapper around such predictors, producing prediction regions $\mathcal{C}(x)$ that have formal guarantees of the form $\mathcal{P}_{X,Y}(Y\notin\mathcal{C}(X))\leq\alpha$ for some prespecified level $\alpha$.
@@
 Generalizations of quantiles have a long history in statistics [30, 31].
-Unlike univariate data, multivariate data does not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering in higher dimensional spaces.
-To account for this, one class of multivariate quantiles, known as “direction quantiles,” prescribes directions along which data points can be compared [32, 33, 34].
-That is, points $x\in\mathbb{R}^{n}$ are projected along some direction $u\in\mathcal{S}^{n-1}$, after which the standard notion of a quantile can be leveraged.
-The notion of the $\alpha$ quantile for a random variable $X\in\mathcal{X}$ clearly then depends on this choice of direction $u$, which we denote by $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$.
+Unlike univariate data, multivariate data do not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering in higher dimensional spaces.
+The notion of a “directional quantile” for a random variable $X\in\mathbb{R}^{n}$ can, however, be directly defined given some direction $u\in\mathcal{S}^{n-1}$, namely as $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$ [32, 33, 34].
 When there is no ambiguity, we just denote it as $Q(\alpha,u)$.
-In cases where domain knowledge assigns an unambiguous notion of multivariate ordering, such knowledge can be encoded into the choice of $u$.
-In other cases, however, a more agnostic approach is employed: rather than choosing any particular direction, all directions are considered simultaneously.
 For any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$.
-To consider all directions simultaneously, the quantile envelope is then defined to be
+The quantile envelope is then the intersection thereof:
 $D(\alpha)=\bigcap_{u\in\mathcal{S}^{n-1}}H(u,Q(\alpha,u)).$ (1)
@@
 2.3 Predict-Then-Optimize
-Prediction regions produced from conformalized predictors are often leveraged differently across classification and regression settings.
-In the case of classification, “regions” simply constitute a subset of the label space, making their direct use by end users straightforward, as demonstrated in works such as [35].
+In the case of classification, conformal prediction regions simply constitute a subset of the label space, making their direct use by end users straightforward [35].
 In high-dimensional regression settings, however, prediction regions become harder to use directly; for this reason, recent works have started shifting focus to using them in their implicit forms.
@@
 The final result is then a decision made with this specification, $w^{*}:=\min_{w}f(w,\widehat{c})$.
-An example of such a setting is if an optimal labor allocation $w^{*}$ is sought based on predicted demand $\widehat{c}$ from observed transactions $x$ in a food delivery platform.
-While the predicted $\widehat{c}$ is often simply trusted in the downstream optimization, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to suboptimal decision-making.
-For this reason, recent interest has been placed on studying a “robust” alternative to this formulation [37, 38, 39].
+An example of such a setting is if an optimal labor allocation $w^{*}$ is sought based on predicted demand $\widehat{c}$ from transactions $x$ in a delivery platform.
+While the predicted $\widehat{c}$ is often trusted, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to suboptimal decision-making.
+For this reason, recent interest has been placed on studying a “robust” formulation [37, 38, 39].
 Following this line of work, [36] proposed studying $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$.
-The primary consequence of doing so was having a probabilistic guarantee on the suboptimality gap.
-That is, denoting the suboptimality gap by $\Delta(x,c):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})-\min_{w}f(w,c)$, where $c$ is the true parameter corresponding to $x$, they established that leveraging conformal uncertainty regions guarantees $\mathcal{P}_{X,C}\left(0\leq\Delta(X,C)\leq L\mathrm{\ diam}(\mathcal{C}(X))\right)\geq 1-\alpha$ if $f(w,c)$ is an objective function that is convex-concave and $L$-Lipschitz in $c$ for any fixed $w$.
 2.4 Related Works
@@
 That is, for an ensembling algorithm $\mathcal{F}:\mathcal{Y}^{K}\rightarrow\mathcal{Y}$, a score function $s(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y)$ would be defined.
-Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ would then be marginally calibrated.
+Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ would then be calibrated.
 Such an approach, however, lacks some desirable properties.
@@
 Generally, however, methods leveraging the full collection of predictors produce less conservative regions [25, 29].
-Such works aggregate the individual prediction regions into a final region by defining
-$\mathcal{C}(x):=\left\{y\mid\sum_{k=1}^{K}w_{k}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]\geq\widehat{a}\right\}$ (2)
-for weights $\{w_{k}\}\in[0,1]$ such that $\sum_{k=1}^{K}w_{k}=1$ and a threshold $\widehat{a}$.
-Methods then differ in the procedure by which $\{w_{k}\}$ and $\widehat{a}$ are prescribed, several of which were prescribed in [29], whose detailed presentation is deferred to Appendix F for space reasons.
-We note that the methods of [25] are designed for a different setting than that considered herein, namely that in which conformal coverage is sought adaptively over online data streams.
-In this vein, [40] is a recent work that has similarly proposed a vector-score extension as that discussed herein, in which candidate weight vectors $\{w_{m}\}\in\mathbb{R}^{K}$ are searched over for score aggregation.
+Such works aggregate the individual prediction regions into a final region by defining $\mathcal{C}(x):=\{y\mid\sum_{k=1}^{K}w_{k}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]\geq\widehat{a}\}$ for weights $\{w_{k}\}\in[0,1]$ such that $\sum_{k=1}^{K}w_{k}=1$ and a threshold $\widehat{a}$.
+Methods then differ in the procedure by which $\{w_{k}\}$ and $\widehat{a}$ are prescribed, several of which were prescribed by [29], whose detailed presentation is deferred to Appendix L for space reasons.
+We note that the methods of [25] are designed for a different setting than that considered herein, namely that in which conformal coverage is sought adaptively over data streams.
+In this vein, [40] have recently proposed a vector-score extension as that discussed herein, in which candidate weight vectors $\{w_{m}\}\in\mathbb{R}^{K}$ are searched over for score aggregation.
 That is, a vector $s(x):=(s_{1}(x,y),...,s_{K}(x,y))\in\mathbb{R}^{K}$ of scores $s_{k}(x,y)$ corresponding to each predictor $f_{k}(x)$ is predicted and its aggregate prediction region defined on the projection $\langle w_{m^{*}},s\rangle$ for $w_{m^{*}}$ the weight resulting in the smallest prediction region.
@@
 3 Method
-We now introduce a procedure for conformal score aggregation for ensembles.
-We discuss the approach and its coverage guarantees in Section 3.1.
-We then discuss the downstream use of its prediction regions in Section 3.2.
 3.1 Multivariate Score Quantile
-We wish to consider the setting typical of conformal model aggregation, as discussed in Section 2.4, in which predictors $f_{1}(x),...,f_{K}(x)$ and corresponding scores $s_{1}(x,y),...,s_{K}(x,y)$ are defined.
-We further assume a similar premise as that taken by [40], in which the individual scores are stacked into a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$.
-A naive approach would then involve leveraging standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$.
+We consider the setting typical of conformal model aggregation, as discussed in Section 2.4, in which predictors $f_{1}(x),...,f_{K}(x)$ and corresponding scores $s_{1}(x,y),...,s_{K}(x,y)$ are defined.
+We assume a similar premise as [40], in which the scores are stacked into a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$.
+A naive approach would then leverage standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$.
 Similar to the naive conformalization of an ensembled predictor discussed in Section 2.4, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties.
@@
 Such a $\widehat{\mathcal{Q}}$ naturally generalizes the standard scalar acceptance interval of $[0,\widehat{q}]$ in the case of non-negative score functions.
-We briefly highlight the distinction between acceptance regions and prediction regions: the former are subsets of the space of multivariate scores, i.e.
-$\subset\mathbb{R}^{K}$, that ultimately define the criteria for retaining particular $y$ values in the prediction region whereas the latter are the subsets of the output space that ultimately have coverage guarantees, i.e.
-$\subset\mathcal{Y}$.
+We briefly highlight the distinction between acceptance regions and prediction regions.
+The former are subsets of the space $\subset\mathbb{R}^{K}$ of multivariate scores that ultimately define the criteria for retaining particular $y$ values in the prediction region.
+The latter are the subsets of the output space $\mathcal{Y}$ and it is these that ultimately have coverage guarantees.
 The two, however, are directly related; in particular, for a fixed score $s(x,y)$, a larger acceptance region will result in a more conservative prediction region.
@@
 The shape of this outer frontier remains fixed over the family and is merely scaled outward from the origin with $t$.
-Under this choice, performing a comparison of $s_{1},s_{2}\in\mathbb{R}^{K}$ in the aforementioned fashion, i.e.
+Under this choice, comparing $s_{1},s_{2}\in\mathbb{R}^{K}$, i.e.
 checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the smallest $t$ for which the outer frontier of $\mathcal{A}_{t}$ intersects $s$.
@@
 The problem of discovering an optimal pre-ordering can, thus, be equivalently stated as seeking to define the outer frontier of $\mathcal{A}_{t}$ to match that of the tightest $1-\alpha$ convex cover of $\mathcal{S}_{C}$.
-This final reframing, therefore, naturally motivates selecting the outer frontier to be the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$.
-Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangeability of its points with future test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$.
-The goal, then, follows as seeking to define the outer frontier as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$.
-For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute the final $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$.
+This motivates selecting the outer frontier to be the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$.
+Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangeability of its points with test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$.
+The goal follows as seeking to define the outer frontier as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$.
+For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$.
 Such a split is predicated on the assumption that the $1-\alpha$ quantile envelope defined over $\mathcal{S}_{C}^{(1)}$ resembles that of $\mathcal{S}_{C}^{(2)}$, implying the $|\mathcal{S}_{C}^{(1)}|$ should be sufficiently large as to capture this structure accurately.
 We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$ using a technique paralleling that described in Section 2.2.
-In particular, we start by selecting the projection directions $\{u_{m}\}$ of Equation 1.
-The restriction of $s$ to the positive orthant induces a natural modification of Equation 1, namely where $u\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$ instead of $\mathcal{S}^{K-1}$.
-To best approximate Equation 1, we wish for $\{u_{m}\}$ to be uniformly distributed over $\mathcal{S}_{+}^{K-1}$; however, exactly finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [41] If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values of $v$.
+In particular, we start by selecting the projection directions $\{u_{m}\}$ of Equation 1; since $s\in\mathbb{R}_{+}^{K}$, we similarly restrict $u_{m}\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$.
+To best approximate Equation 1, we wish for $\{u_{m}\}$ to be uniformly distributed over $\mathcal{S}_{+}^{K-1}$; however, exactly finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [41] If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values.
 We now wish to define the quantile thresholds $\{\widetilde{q}_{m}\}$ for the selected directions to optimally capture $1-\alpha$ of $\mathcal{S}_{C}^{(1)}$.
@@
 If the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$.
-Importantly, while the aforementioned procedure will necessarily result in convex regions $\widehat{\mathcal{Q}}$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, as discussed in Section 3.2.
-Such flexibility is necessary, as many studies have demonstrated the need for nonconvex prediction regions for downstream utility, such as [36, 42, 43].
+Critically, such calculations can be computed efficiently in vector form.
+Due to space restrictions, we defer this discussion to Appendix D.
+We additionally there empirically validate the efficiency of the procedure under this vectorized implementation.
+We present the full algorithm in Algorithm 1.
+Importantly, while this procedure will result in convex regions $\widehat{\mathcal{Q}}$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, as discussed in Section 3.2.
 However, it is unsurprising such flexibility exists, as even a single scalar score $s_{1}(x,y)$ can produce nonconvex prediction regions.
-We present the full algorithm in Algorithm 1.
+One additional notable property of the CSA prediction regions is that their sizes vary across $x$ even if such variability is not baked into the constituent scores.
+For instance, using $s_{k}(x,y):=|f_{k}(x)-y|$ with standard, scalar conformal prediction yields intervals of length $2\widehat{q}_{k}$ for any $x$, yet $|\mathcal{C}^{\mathrm{CSA}}(x)|$ even with such $\{s_{k}(x,y)\}$ will vary with $x$.
+This variability is desirable, as predictive uncertainty is seldom uniform across the covariate space.
+See Appendix E for a full illustration of this.
 Algorithm 1 CSA: $\textsc{UnifHypersphere}(K)$ is an assumed subroutine that samples $\sim\mathrm{Unif}(\mathcal{S}^{K-1})$.
@@
 2:  $[\beta_{\mathrm{lo}},\beta_{\mathrm{hi}}]\leftarrow[\alpha/M,\alpha],\widehat{\mathcal{Q}}\leftarrow\emptyset$
-3:  $\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{k}(x_{i},y_{i}))_{k=1}^{K}\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}}$
-4:  $\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
-5:  while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
-6:     $\beta\leftarrow\frac{\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}}}{2}$
-7:     $\left\{\widetilde{q}_{m}\leftarrow(1-\beta)\text{ emp. quantile: }\{u_{m}^{\top}s_{i}\}_{s_{i}\in S_{C}^{(1)}}\right\}_{m=1}^{M}$
-8:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
-9:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$
-10:     else $\beta_{\mathrm{hi}}\leftarrow\beta$
-11:  end while
-12:  $\widehat{t}\leftarrow(1-\alpha)\text{ emp. quantile: }\{\max_{m\in[M]}\frac{u_{m}^{\top}s_{i}}{\widetilde{q}_{m}}\}_{s_{i}\in S_{C}^{(2)}}$
-13:  Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
+3:  $\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{k}(x_{i},y_{i}))_{k=1}^{K}\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}},\quad\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
+4:  while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
+5:     $\beta\leftarrow(\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}})/2$
+6:     $\left\{\widetilde{q}_{m}\leftarrow(1-\beta)\text{ empirical quantile of }\{u_{m}^{\top}s_{i}\}_{s_{i}\in S_{C}^{(1)}}\right\}_{m=1}^{M}$
+7:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
+8:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$ else $\beta_{\mathrm{hi}}\leftarrow\beta$
+9:  end while
+10:  $\widehat{t}\leftarrow(1-\alpha)\text{ empirical quantile of }\{\max_{m\in[M]}(u_{m}^{\top}s_{i}/\widetilde{q}_{m})\}_{s_{i}\in S_{C}^{(2)}}$
+11:  Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
 3.2 Predict-Then-Optimize
 With this generalization of the score function, a natural question is how to leverage the resulting prediction regions $\mathcal{C}(x)$.
-For both classification and regression, $\mathcal{C}(x)=\{y\mid s(x,y)\in\widehat{\mathcal{Q}}\}$, or $\mathcal{C}(x)=\bigcap_{m=1}^{M}\mathcal{C}_{m}(x)$ where $\mathcal{C}_{m}(x):=\{y\mid u_{m}^{\top}s(x,y)\leq\widehat{q}_{m}\}$.
-In the classification setting, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward.
-For a fixed $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the thresholds $\widehat{q}_{m}$ after projection.
+For both classification and regression, $\mathcal{C}(x)=\bigcap_{m=1}^{M}\mathcal{C}_{m}(x)$ where $\mathcal{C}_{m}(x):=\{y\mid u_{m}^{\top}s(x,y)\leq\widehat{q}_{m}\}$.
+For classification, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward: for any $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the thresholds $\widehat{q}_{m}$ after projection.
 In the case of regression, however, the prediction region cannot be explicitly constructed in the general case, since $\mathcal{Y}$ contains uncountably many elements.
 In fact, explicit construction is generally not of interest for downstream regression applications.
-Instead, prediction regions in regression settings are often simply used in their implicit form.
 We, therefore, focus on one particular application, namely that of [36] discussed in Section 2.3, and demonstrate the CSA prediction regions can be leveraged in their framework.
@@
 The GPCP score, used to define the score components, is
-$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}\left[d_{k}\left(\widehat{c}_{kj},c\right)\right],$ (3)
-where $d_{k}(\widehat{c},c)$ is a norm over the $C$ space.
+$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}\left[||\widehat{c}_{kj}-c||_{2}\right].$ (2)
 Notably, this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X)=\delta(f_{k}(X))$.
@@
 Notably then, the projection $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)$ is convex in $c$, since the projection directions are all restricted to $\mathcal{S}_{+}^{K-1}$.
-As a result, we can compute the maximizer $c^{*}$, since
-$c_{\vec{j}}^{*}$ $:=\operatorname*{arg\,max}_{c}f(w,c)$ (4)
-   s.t. $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\quad\forall m\in\{1,...,M\}$
+Thus,
+[…]
 remains a standard convex optimization problem.
@@
 This coupled with the fact that computing over these indices is trivially parallelizable means this approach is still computationally tractable.
-The full procedure is outlined in Algorithm 2.
-Algorithm 2 Predict-Then-Optimize Under CSA
-1:  Inputs: Context $x$, Predictors $\{q_{k}(C\mid X)\}_{k=1}^{K}$, Optimization steps $T$, Sample counts $\{J_{k}\}_{k=1}^{K}$, CSA quantile $\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$
-2:  $\{\{\widehat{c}_{kj}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X)\}_{k=1}^{K},\mathcal{J}=\prod_{k=1}^{K}[J_{k}]$
-3:  $w^{(0)}\sim U(\mathcal{W})$
-4:  for $t\in\{1,\ldots T\}$ do
-5:     for $\vec{j}\in\mathcal{J}$ do $c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}f(w^{(t)},c)$$\mathrm{s.t.}\quad\forall m\in 1,...,M\qquad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}$
-6:     $c^{*}\leftarrow\operatorname*{arg\,max}_{c_{\vec{j}}^{*}}f(w^{(t)},c_{\vec{j}}^{*})$
-7:     $w^{(t)}\leftarrow\Pi_{\mathcal{W}}(w^{(t-1)}-\eta\nabla_{w}f(w^{(t-1)},c^{*}))$
-8:  end for
-9:  Return $w^{(T)}$
-Table 1: Average coverages across tasks for $\alpha=0.10$ are shown in the top rows and average prediction set sizes in the bottom rows.
+The full procedure is outlined in Algorithm 2 in Section C.1 due to space limitations.
+4 Experiments
+We now study CSA empirically across several tasks, demonstrating its coverage guarantees with reduced conservatism.
+We demonstrate improvements in an ImageNet classification task in Section 4.1, across real-data regression benchmark tasks in Section 4.2 as proposed by [42], and in a downstream predict-then-optimize task in Section 4.3.
+We additionally assess the robustness of CSA to imbalanced ensembles and perform an ablation study of the two-stage calibration.
+We note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment, meaning care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for CSA, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction.
+We note the splits in each of the sections that follow.
+We compare against the methods presented in Section 2.4, viz.
+the model selection of [27], the aggregation methods of [25], and the single weighted score projection (VFCP) of [40].
+We additionally include the initial strategy discussed in Section 2.4, in which the ensemble predictor is directly conformalized, using a natural aggregate “ensemble” score, given in the following sections.
+From the work of [25], we consider the following methods: the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches (see Appendix L).
+Notably, these methods do not lend themselves for use in the predict-then-optimize setting, so we eliminate them from consideration therein.
+VFCP can only be applied in classification settings; we, thus, do not compare to it across the regression tasks.
+Code will be made public upon acceptance.
+4.1 Classification Tasks
+Table 1: Classification results are shown across tasks for $\alpha=0.10$, $\alpha=0.05$, and $\alpha=0.01$, with coverages in the top (grey) and average prediction set sizes (white) in the bottom of each row.
 Both were assessed over a batch of i.i.d.
@@
 Standard deviations and means were computed across 10 randomized draws of the calibration and test sets.
-Dataset Models ResNet VGG DenseNet VFCP $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA
-ImageNet Coverage 0.901 (0.005) 0.902 (0.003) 0.902 (0.003) 0.899 (0.004) 0.938 (0.003) 0.909 (0.004) 0.9 (0.004) 0.9 (0.003)
-Size 137.004 (1.98) 136.116 (2.206) 120.096 (2.427) 46.063 (1.089) 87.337 (1.604) 82.746 (1.692) 131.856 (2.378) 34.006 (0.924)
-4 Experiments
-We now study CSA empirically across several tasks, demonstrating its coverage guarantees with reduced conservatism.
-We demonstrate improvements in an ImageNet classification task in Section 4.1, in real-data UCI regression tasks as studied in [29] in Section 4.2, and in a downstream predict-then-optimize task in Section 4.3.
-To allow for fair comparison, we note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment.
-This in turn means that some care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for CSA, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction.
-For this reason, we fixed the split to be $20\%$-$80\%$ in experiments.
-We herein compare against the methods presented in Section 2.4, namely the model selection of [27], the aggregation methods of [25], and the single weighted score projection (VFCP) of [40].
-We specifically consider the following aggregation methods from [25]: the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches, fully described in Appendix F.
-Notably, these methods do not lend themselves for use in the predict-then-optimize setting, so we eliminate them from consideration therein.
-Similarly, VFCP can only be applied in classification settings; we, thus, do not compare to it across the regression tasks.
-Code will be made public upon acceptance.
-4.1 Classification Tasks
+Dataset/$\boldsymbol{\alpha}$ ResNet VGG DenseNet VFCP $\boldsymbol{\mathcal{C}^{M}}$ $\boldsymbol{\mathcal{C}^{R}}$ $\boldsymbol{\mathcal{C}^{U}}$ Ensemble CSA
+ImageNet 0.901 (0.005) 0.902 (0.003) 0.902 (0.003) 0.899 (0.004) 0.938 (0.003) 0.909 (0.004) 0.9 (0.004) 0.899 (0.004) 0.9 (0.003)
+$(\alpha=0.10)$ 137.004 (1.98) 136.116 (2.206) 120.096 (2.427) 46.063 (1.089) 87.337 (1.604) 82.746 (1.692) 131.856 (2.378) 69.123 (1.317) 34.006 (0.924)
+$(\alpha=0.05)$ 0.95 (0.003) 0.949 (0.004) 0.952 (0.002) 0.95 (0.003) 0.975 (0.002) 0.954 (0.004) 0.95 (0.003) 0.949 (0.002) 0.95 (0.003)
+220.022 (2.072) 229.523 (3.076) 208.658 (2.016) 78.108 (2.004) 166.933 (2.157) 143.323 (2.932) 220.491 (2.773) 112.161 (2.115) 59.574 (3.382)
+$(\alpha=0.01)$ 0.99 (0.001) 0.991 (0.001) 0.989 (0.002) 0.99 (0.001) 0.997 (0.001) 0.991 (0.002) 0.99 (0.002) 0.99 (0.002) 0.99 (0.002)
+491.952 (6.353) 726.028 (12.157) 459.399 (6.739) 194.691 (4.579) 580.592 (7.715) 532.155 (24.829) 559.188 (7.07) 299.453 (6.526) 201.32 (46.509)
 We first study the predictive efficiency of the aforementioned methods on the ImageNet classification task [43].
@@
 Conformalization on the individual models was performed using the standard classification score function across all approaches, namely $s(x,y)=\sum_{j=1}^{l}\widehat{f}(x)_{\pi_{j}(x)}$ where $y=\pi_{l}(x)$ and $\pi(x)$ is the permutation of $\{1,\ldots,|\mathcal{Y}|\}$ that sorts $\widehat{f}\left(x\right)$ from most to least likely.
+Here, the “Ensemble” score was computed with the same $s(x,y)$, replacing $f_{k}(x)$ with the ensemble average probability, i.e.
+$\mu(x)_{j}:=\sum_{k}(f_{k}(x))_{j}/K$.
 Calibration was performed using 85% of the ImageNet test set and assessment of the coverage and interval lengths on the remaining 15%, with 10 trials conducted over randomized draws of these calibration and test sets.
-The results are presented in Table 1.
-We see that all the approaches exhibit the desired coverage guarantees.
-However, the CSA score approach consistently produce significantly smaller prediction regions than both the individually conformalized input views and alternate aggregation strategies.
-4.2 UCI Regression Tasks
-We now similarly study the predictive efficiency of CSA across a suite of regression tasks from the UCI repository [45] as studied by [29].
-We specifically consider those tasks for which there are at least 1,000 samples; the complete list of tasks are provided in Appendix D.
-The data for each task were split with 80% for training, 10% for calibration, and 10% for testing coverage and interval lengths.
-The problem setup was replicated from [29], in which four prediction methods were ensembled for each prediction task, namely an OLS model, a LASSO linear model, a random forest, and a multi-layer perceptron.
-A standard absolute value residual function was used as the score across all methods, namely $s(x,y)=|\widehat{f}(x)-y|$.
-Prediction intervals could be explicitly constructed for the $\mathcal{C}^{M}$, $\mathcal{C}^{R}$, and $\mathcal{C}^{U}$ aggregation methods.
-For assessing interval length for CSA, however, a discretized grid $\mathcal{G}_{Y}\subset\mathcal{Y}$ of coarseness $\Delta y$ was considered, and an interval length estimate given by $\mathcal{L}(\mathcal{C}(x))\approx\Delta y\cdot|\{y:y\in\mathcal{G}_{Y},s(x,y)\in\widehat{\mathcal{Q}}\}|$.
-$10$ trials were conducted for each task, randomized over subset selections of the training, calibration, and test sets.
-We also present an ablation, labeled “CSA (Single-Stage),” to demonstrate the two-stage calibration described in Section 3.1 is necessary to retain coverage; this single-stage approach does not split $\mathcal{S}_{C}$ and instead directly computes $\{\widehat{q}_{m}\}$ on $\mathcal{S}_{C}$ per Section 3.1.1.
-A subset of the results are given in Table 2; the full set of results is deferred to Appendix D due to space limitations.
-As in the results presented in Section 4.1, we see that CSA retains the coverage guarantees typical of conformal prediction yet produces significantly smaller prediction intervals than both the individual models and the alternate aggregation strategies.
-We additionally see that the “CSA (Single-Stage)” approach fails to retain coverage, demonstrating the necessity of the two-stage calibration.
-Table 2: Average coverages across tasks for $\alpha=0.05$ are shown in the top rows and average prediction set lengths in the bottom rows, both were assessed over a batch of i.i.d.
-test samples (10% of the datasets).
-Standard deviations and means were computed across 10 randomized draws of the training, calibration, and test sets.
-Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, thus, precluded from comparison.
-Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
-3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
-Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
-parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
-Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
-pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
-Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
-pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
-Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
-slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
-Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
+A $25/75\%$ split was used for $\mathcal{D_{C}}^{(1)}$-$\mathcal{D_{C}}^{(2)}$.
+The results are presented in Table 1; the full results across additional $\alpha$ is given in Appendix F.
+We see that all the approaches exhibit the desired coverages across $\alpha$.
+However, CSA consistently produces significantly smaller prediction regions than both the individually conformalized models and alternate aggregation strategies.
+4.2 Regression Tasks
+We now similarly study the predictive efficiency of CSA across a suite of regression tasks from [42].
+The data for each task were split with 50/45/5% for training, calibration, and testing for coverage and interval lengths, with five trials conducted over randomized selections of such sets.
+A $5/95\%$ split was used for $\mathcal{D_{C}}^{(1)}$-$\mathcal{D_{C}}^{(2)}$.
+The problem setup was replicated from [29], in which four prediction methods were ensembled, namely an OLS model, a LASSO linear model, a random forest (RF), and an XGBoost model.
+A residual function was used as the score across all methods, namely $s(x,y)=|\widehat{f}(x)-y|$.
+Here, the “Ensemble” score was the standard $s(x,y):=\frac{|\mu(x)-y|}{\sigma(x)}$, where $(\mu(x),\sigma(x))$ are the ensemble mean and standard deviation.
+Prediction intervals could be analytically constructed for the $\mathcal{C}^{M}$, $\mathcal{C}^{R}$, and $\mathcal{C}^{U}$ methods.
+To assess CSA, however, a discretized grid $\mathcal{G}_{Y}\subset\mathcal{Y}$ of coarseness $\Delta y$ was considered, and an interval length estimate given by $\mathcal{L}(\mathcal{C}(x))\approx\Delta y\cdot|\{y:y\in\mathcal{G}_{Y},s(x,y)\in\widehat{\mathcal{Q}}\}|$.
+We also present an ablation, labeled “Single-Stage,” to demonstrate the two-stage calibration is necessary to retain coverage; this single-stage approach does not split $\mathcal{S}_{C}$ and instead directly computes $\{\widehat{q}_{m}\}$ on $\mathcal{S}_{C}$ per Section 3.1.1.
+We provide the results for $\alpha=0.05$ and $\alpha=0.025$ to demonstrate the consistency of the method performance.
+A subset of the results is given in Table 2; the full set of results is deferred to Appendix G.
+As in the results of Section 4.1, we see that CSA retains the coverage guarantees typical of conformal prediction yet produces significantly smaller prediction intervals than both the individual models and the alternate aggregation strategies.
+We additionally see that the “Single-Stage” approach fails to retain coverage, demonstrating the necessity of the two-stage calibration.
+We provide a visual comparison of the prediction regions resulting from these methods in Appendix H.
+We additionally assessed the robustness of our method to imbalanced ensembles.
+The experiments of [29] were conducted on a UCI benchmark task [44] with an ensemble of an OLS model, a LASSO linear model, a random forest, and an MLP, and they found the conformalized random forest to outperform all the proposed aggregation strategies, due to the lack of orthogonal information in considering the other predictors.
+We find that, in these degenerate cases, where the best decision is to simply choose a single predictor, our method outperforms other aggregation methods and nearly matches the performance of the best conformalized predictor in hindset; the results are presented in Appendix I across a number of UCI benchmarks.
+Table 2: The results for five distinct tasks are shown below for $\alpha=0.05$ (top five rows) and $\alpha=0.025$ (bottom five rows).
+For each, the average coverages (grey rows) and prediction set lengths (white rows) with standard deviations are given, both assessed over 5 randomized draws of the training, calibration, and test sets.
+In cases where the method failed to achieve sufficient coverage (i.e.
+<.$<.93$ for $\alpha=0.05$ and $<0.96$ for $\alpha=0.025$), we do not include it in comparison for set length.
+Dataset/$\boldsymbol{\alpha}$ OLS LASSO RF XGBoost $\boldsymbol{\mathcal{C}^{M}}$ $\boldsymbol{\mathcal{C}^{R}}$ $\boldsymbol{\mathcal{C}^{U}}$ Ensemble Single-Stage CSA
+361234 0.97 (0.011) 0.966 (0.011) 0.939 (0.002) 0.954 (0.006) 0.956 (0.011) 0.948 (0.01) 0.96 (0.013) 0.95 (0.006) 0.955 (0.013) 0.957 (0.01)
+$(\alpha=0.05)$ 9.673 (0.160) 9.645 (0.154) 10.080 (0.160) 9.157 (0.052) 9.196 (0.123) 8.703 (0.086) 9.524 (0.056) 17.759 (0.275) 7.646 (0.073) 7.688 (0.181)
+361235 0.947 (0.0) 0.945 (0.005) 0.968 (0.016) 0.95 (0.005) 0.955 (0.016) 0.897 (0.005) 0.953 (0.011) 0.932 (0.021) 0.745 (0.011) 0.984 (0.005)
+$(\alpha=0.05)$ 20.961 (0.651) 24.241 (0.246) 10.096 (0.587) 11.387 (0.452) 11.782 (0.057) 16.088 (0.118) 15.823 (1.272) 6.162 (0.458) 11.695 (0.266)
+361236 0.975 (0.008) 0.975 (0.008) 0.961 (0.0) 0.948 (0.012) 0.948 (0.012) 0.938 (0.012) 0.965 (0.008) 0.934 (0.004) 0.94 (0.004) 0.963 (0.004)
+[…]
+361237 0.969 (0.023) 0.969 (0.023) 0.981 (0.0) 0.923 (0.0) 0.954 (0.015) 0.9 (0.008) 0.969 (0.023) 0.885 (0.038) 0.8 (0.015) 0.977 (0.008)
+$(\alpha=0.05)$ 44.019 (0.990) 44.069 (1.115) 27.035 (1.014) 26.524 (1.244) 31.967 (1.118) 14.473 (0.503) 23.145 (0.199)
+361241 0.954 (0.001) 0.956 (0.001) 0.944 (0.005) 0.957 (0.002) 0.954 (0.002) 0.923 (0.0) 0.952 (0.0) 0.949 (0.001) 0.917 (0.006) 0.951 (0.001)
+$(\alpha=0.05)$ 19.133 (0.062) 20.245 (0.095) 18.102 (0.055) 18.482 (0.062) 17.958 (0.062) 18.932 (0.034) 29.548 (0.191) 15.199 (0.427) 17.328 (0.097)
+361234 0.987 (0.008) 0.987 (0.008) 0.974 (0.004) 0.977 (0.008) 0.982 (0.008) 0.971 (0.01) 0.981 (0.01) 0.97 (0.008) 0.976 (0.01) 0.973 (0.006)
+$(\alpha=0.025)$ 11.939 (0.137) 11.871 (0.084) 12.484 (0.168) 11.972 (0.009) 11.587 (0.110) 11.157 (0.086) 11.965 (0.050) 25.598 (0.974) 9.306 (0.259) 8.855 (0.059)
+361235 0.987 (0.0) 0.982 (0.011) 0.979 (0.011) 0.984 (0.005) 0.989 (0.005) 0.966 (0.016) 0.976 (0.005) 0.958 (0.021) 0.889 (0.011) 0.989 (0.005)
+$(\alpha=0.025)$ 24.595 (0.825) 28.841 (1.129) 11.811 (0.992) 14.237 (0.786) 14.472 (0.172) 12.278 (0.026) 19.231 (0.356) 7.719 (0.467) 12.563 (0.766)
+361236 0.992 (0.004) 0.992 (0.004) 0.981 (0.0) 0.965 (0.008) 0.975 (0.008) 0.965 (0.008) 0.977 (0.012) 0.955 (0.008) 0.955 (0.012) 0.973 (0.004)
+[…]
+361237 0.981 (0.0) 0.981 (0.0) 0.981 (0.0) 0.977 (0.008) 0.962 (0.0) 0.962 (0.0) 0.977 (0.008) 0.965 (0.008) 0.927 (0.031) 0.981 (0.0)
+$(\alpha=0.025)$ 47.738 (0.542) 47.440 (0.959) 30.785 (0.037) 26.208 (0.897) 30.554 (0.561) 27.182 (0.803) 35.982 (0.619) 67.660 (6.380) 18.214 (0.436) 26.897 (0.515)
+361241 0.979 (0.001) 0.978 (0.001) 0.976 (0.001) 0.978 (0.001) 0.978 (0.0) 0.964 (0.002) 0.977 (0.0) 0.972 (0.002) 0.958 (0.003) 0.979 (0.0)
+$(\alpha=0.025)$ 21.772 (0.085) 23.089 (0.106) 21.543 (0.009) 21.454 (0.109) 20.862 (0.088) 19.291 (0.060) 21.905 (0.041) 40.082 (0.045) 17.765 (0.329) 19.897 (0.062)
 4.3 CSA Predict-Then-Optimize
-We now study the use of CSA for a real-world predict-then-optimize traffic routing task, as studied in the conformal predict-then-optimize paper [36].
-In this task, a time series of $T$ previously observed precipitation readings are used to predict future precipitations and, in turn, anticipated traffic, as fully described in Appendix E.
+We now study a real-world predict-then-optimize traffic routing task, from [36].
+In this task, a time series of $T$ preceding precipitations is used to predict future precipitations and, in turn, future traffic, as fully described in Appendix J.
 We consider the traffic routing problem for a fixed source-target pair $(s,t)$ over the graph of Manhattan, where $|\mathcal{V}|=4584$ and $|\mathcal{E}|=9867$.
-Formally, such a problem seeks
-$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}\widehat{c}^{T}w$ (5)
-[…]
+Formally,
+$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}\widehat{c}^{T}w\quad\textrm{s.t.}\quad w\in[0,1]^{\mathcal{E}},Aw=b,\mathcal{P}_{X,C}(C\in\mathcal{C}(X))\geq 1-\alpha$
 where $x\in\mathbb{R}^{T\times H\times W}$ are the previous precipitation readings, $w_{e}\in\mathbb{R}^{|\mathcal{E}|}$ the traffic proportion routed along road $e$, $c\in\mathbb{R}^{|\mathcal{E}|}$ the transit times anticipated across roads, $A\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{E}|}$ the graph incidence matrix, and $b\in\mathbb{R}^{|\mathcal{V}|}$ the vector that specifies the routing problem, in which $b_{s}=1,b_{t}=-1,$ and $b_{k}=0$ for $s$ the travel source node, $t$ the terminal node, and $k\notin\{s,t\}$ all other nodes.
 We then consider two probabilistic models for traffic prediction, namely one based on the classical probabilistic Lagrangian integro-difference approach (STEPS) of [45] and one on the modern latent diffusion model (LDM) approach of [46].
-As a result of the higher inference cost of the latter, we consider the setup where $J_{1}>J_{2}$, specifically with $J_{1}=4$ and $J_{2}=1$, which additionally serves to empirically highlight the flexibility of non-uniform sampling from predictors discussed in Section 3.2.
+As a result of the higher inference cost of the latter, we consider the setup where $J_{1}>J_{2}$, specifically with $J_{1}=4$ and $J_{2}=1$, highlighting the flexibility of non-uniform sampling from predictors discussed in Section 3.2.
 As discussed in Section 4, the alternate aggregation strategies do not lend themselves for use in this setting.
-We, therefore, only compare CSA to the best model selection strategy, specifically by comparing to the separate conformalizations of the two constituent predictors.
-Each model was conformalized with Equation 3 taken to be the score, where $d_{k}$ was set to be the standard 2-norm across both predictors, such that $s_{1}(x,c)=\min_{j\in 1,...,J_{1}}||\widehat{c}_{1j}-c||_{2}$ for $q_{1}(C\mid x)$ and similarly for $q_{2}(C\mid x)$.
+We, therefore, only compare CSA to the separate conformalizations of the two predictors, with the score from Equation 2.
 We here evaluate the methods using the expected suboptimality gap proportion, $\Delta_{\%}=\mathbb{E}_{X}[\Delta(X,C(X))/\min_{w}f(w,C(X))],$ where $\Delta$ is defined as discussed in Section 2.3.
-This quantity measures the conservatism of the produced robust optimal value and is bounded in $[0,1]$ in this setting.
-Experiments were conducted with $|\mathcal{D}_{C}|=200$.
+This measures the conservatism of the robust optimal value and is bounded in $[0,1]$.
+Experiments were conducted with $|\mathcal{D}_{C}|=200$, with a $20/80\%$ split used for $\mathcal{D_{C}}^{(1)}$-$\mathcal{D_{C}}^{(2)}$.
 The suboptimality was then computed across 100 i.i.d.
 test samples.
-To, therefore, assess the improvement offered by CSA, we conduct two paired t-tests, with hypotheses
-$H_{0}:\Delta^{(\mathrm{CSA})}_{\%}=\Delta^{(\mathrm{STEPS})}_{\%}\quad H_{1}:\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$
-and similarly for $\Delta^{(\mathrm{CSA})}_{\%}$ and $\Delta^{(\mathrm{LDM})}_{\%}$.
-The coverages and results of such t-tests are provided in Table 3, from which we find that CSA significantly reduces the suboptimality after accounting for Bonferroni multiple testing.
+To assess the improvement, we conducted two paired t-tests, where $H_{0}:\Delta^{(\mathrm{CSA})}_{\%}=\Delta^{(\mathrm{STEPS})}_{\%}$ and $H_{1}:\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$ and similarly for $\Delta^{(\mathrm{CSA})}_{\%}$ and $\Delta^{(\mathrm{LDM})}_{\%}$.
+The results are provided in Table 3, from which we find that CSA significantly reduces the suboptimality after accounting for Bonferroni multiple testing.
 We see that, while conformalization of either of the two views individually already produces the desired coverage, CSA produces more informative prediction regions, and hence less conservative robust upper bounds.
-Table 3: Coverages across tasks for $\alpha=0.05$ are shown in the left table across both the individually conformalized and CSA fused approaches.
-P-values of the paired t-tests comparing gap proportions ($\Delta_{\%}$) are shown in the right table.
-Both were computed over 100 i.i.d.
+Table 3: Coverages for $\alpha=0.05$ for the individually conformalized and CSA approach and p-values of the paired t-tests comparing $\Delta_{\%}$ are shown, both computed over 100 i.i.d.
 test samples.
-STEPS LDM CSA
-0.981 0.962 0.968
-$H_{1}$ p-value
-$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$ $3.61\times 10^{-4}$
-$\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{LDM})}_{\%}$ $9.50\times 10^{-4}$
+Coverage P-values for $H_{1}$
+STEPS LDM CSA $\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{STEPS})}_{\%}$: $3.61\times 10^{-4}$
+0.981 0.962 0.968 $\Delta^{(\mathrm{CSA})}_{\%}<\Delta^{(\mathrm{LDM})}_{\%}$: $9.50\times 10^{-4}$
 5 Discussion
-We have presented CSA, a framework for producing informative prediction regions in ensemble predictor pipelines.
-We additionally demonstrated the generality of the proposed pipeline in settings of both classification and robust predict-then-optimize regression tasks.
-This work suggests many directions for extension.
-One point of interest is the integration of CSA with other downstream regression applications.
-For instance, [48] proposed an end-to-end differentiable extension to [36]: extending CSA for integration to their framework would be of great interest.
-Additionally, given the prevalence of late-stage fusion in robotics applications, an exciting applied avenue of research would investigate the use of CSA in settings of robust control.
+We have presented a framework for producing informative prediction regions in ensemble predictor pipelines, suggesting many directions for extension.
+One is in extracting insights of the relative predictor uncertainties from the data-driven $\lesssim$.
+Another is the integration of CSA with [47], which proposed an end-to-end extension to [36].
+Finally, given the prevalence of sensor fusion in robotics, another avenue is to study the use of CSA in robust control.
 Appendix A CSA Visual Walkthrough
@@
 Note that the construction of this feasible is never explicitly done in practice and is only implicitly used by convex solver routines in practice.
+We can, therefore, then solve Equation 10 over all possible $\vec{j}\in\mathcal{J}$ and aggregate the maxima to compute $c^{*}$.
 Figure 8: A candidate $c\in\mathcal{C}$ is in prediction region if the projections of its distances $(d_{1},d_{2})$ from at least one pair of points indexed by $\vec{j}:=(j_{1},j_{2})$ is in the acceptance region $\widehat{\mathcal{Q}}$.
@@
 Note that the left is again a visualization over the $\mathcal{C}$ space, whereas the right is of the score space.
-We can, therefore, then solve Equation 12 over all possible $\vec{j}\in\mathcal{J}$ and aggregate the maxima to compute $c^{*}$.
-Appendix D Additional UCI Results
-We consider those regression tasks from the UCI repository [45] that have at least 1,000 samples. The complete list of tasks considered herein is provided in Table 4.
-Table 4: Tasks considered for empirical validation of prediction region efficiency and coverage, along with their relevant dimensions. All tasks are scalar regression tasks, i.e. $y\in\mathbb{R}$.
-Dataset Observations Input Dimension
-3droad 434874 3
-airfoil 1503 5
-bike 17379 17
-parkinsons 5875 20
-pol 15000 26
-pumadyn32nm 8192 32
-slice 53500 385
+C.1 CSA Predict-Then-Optimize Algorithm
+We provide the condensed presentation of the predict-then-optimize algorithm below.
+Algorithm 2 Predict-Then-Optimize Under CSA
+1:  Inputs: Context $x$, Predictors $\{q_{k}(C\mid X)\}_{k=1}^{K}$, Optimization steps $T$, Sample counts $\{J_{k}\}_{k=1}^{K}$, CSA quantile $\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$
+2:  $\{\{\widehat{c}_{kj}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X)\}_{k=1}^{K},\mathcal{J}=\prod_{k=1}^{K}[J_{k}]$
+3:  $w^{(0)}\sim U(\mathcal{W})$
+4:  for $t\in\{1,\ldots T\}$ do
+5:     for $\vec{j}\in\mathcal{J}$ do $c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}f(w^{(t)},c)\qquad\mathrm{s.t.}\quad\forall m\in 1,...,M\qquad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}$
+6:     $c^{*}\leftarrow\operatorname*{arg\,max}_{c_{\vec{j}}^{*}}f(w^{(t)},c_{\vec{j}}^{*})$
+7:     $w^{(t)}\leftarrow\Pi_{\mathcal{W}}(w^{(t-1)}-\eta\nabla_{w}f(w^{(t-1)},c^{*}))$
+8:  end for
+9:  Return $w^{(T)}$
+Appendix D Computational Efficiency
+D.1 Vectorized Score Computation
+We now discuss the vectorized form of the computations discussed in Section 3.1.2.
+In particular, putting the scores into a matrix $S_{C}^{(1)}=[s_{1},...,s_{N_{C_{1}}}]^{\top}\in\mathbb{R}^{N_{C_{1}}\times K}$ and directions into a matrix $U=[u_{1},...,u_{M}]^{\top}\in\mathbb{R}^{K\times M}$, all the projections $u_{m}^{\top}s_{i}$ can be computed as $S_{C}^{(1)}U\in\mathbb{R}^{N_{C_{1}}\times M}$, where $[S_{C}^{(1)}U]_{i,m}$ is precisely $u_{m}^{\top}s_{i}$.
+$\tilde{q}\in\mathbb{R}^{M}:=\{\tilde{q}_{m}:=\mathrm{quantile}([S_{C}^{(1)}U]_{:,m};1-\alpha)\}$ is then the quantile per row.
+For any test point $s\textquoteright\in\mathbb{R}^{K}$, we can then very efficiently check if it falls into the region by checking if it satisfies $Us\textquoteright\leq\tilde{q}$ component-wise.
+Each iteration of the loop to find $\beta^{*}$, therefore, is very fast, and we find the search typically converges in 5-10 iterations.
+The final step is computing $\widehat{t}$.
+Computing this follows similarly to above, where we take the scores $S_{C}^{(2)}$, compute projections $S_{C}^{(2)}U\in\mathbb{R}^{N_{C_{2}}\times M}$, find $\widetilde{T}:=S_{C}^{(2)}U/\tilde{q}\in\mathbb{R}^{N_{C_{2}}\times M}$, where division is interpreted as being defined component-wise along the rows, computing the maxima similarly along the rows $[T^{*}]_{i}:=\max\widetilde{T}_{i,:}$, and finally computing $\widehat{t}$ as the $1-\alpha$ quantile of $T^{*}$.
+D.2 Empirical Efficiency Validation
+To demonstrate the computational efficiency of this vectorized approach, we reran the experiment on the “Parkinsons” UCI task with both varying numbers of predictors ($K$) and projection directions ($M$); for each combination, we measured the total time taken to compute the quantile (i.e.
+to run Algorithm 1) and to perform the projection to assess coverage for the test points.
+The additional predictors were taken to be random forests with different numbers of trees.
+$K$ is given in the left column and $M$ in each column heading, with the entry for each $(K,M)$ pair being reported in seconds.
+As expected, by the vectorized nature of the computations, as discussed in the main paper, the performance scales gracefully over $M$ at roughly $\mathcal{O}(M)$ and remains roughly constant in $K$.
+Table 4: Performance values for varying $K$ (number of predictors) and $M$ (number of projection directions). All values are reported in seconds.
+$K$ $M$ 10 100 1000 10000
+6 0.111668 0.373029 2.32803 37.2561
+8 0.0961056 0.327051 2.16211 36.7216
+10 0.117146 0.373464 2.73875 37.2603
+12 0.123772 0.384527 2.35386 37.0735
+Appendix E Prediction Region Intuition
+We now consider a simplified setting of the general procedure to gain insight into the efficiency resulting prediction regions.
+In particular, we consider a scalar regression setting with $K$ predictors $f_{1},...,f_{K}:\mathcal{X}\rightarrow\mathbb{R}$.
+We further suppose, with a slight abuse of notation, $f(x)-y\sim\mathcal{N}(0,\Sigma)$ for $f(x):=[f_{1}(x),...,f_{K}(x)]$ and that the scores are taken to be $s_{k}(x,y):=(f_{k}(x)-y)^{2}$.
+Then, as $\widehat{\mathcal{Q}}$ is precisely the region with minimal volume that captures $1-\alpha$ density, it is precisely given by $\chi_{K,1-\alpha}^{2}$, the $1-\alpha$ quantile of the $\chi^{2}$ distribution with $K$ degrees of freedom.
+That is, the prediction regions are $\mathcal{C}^{\mathrm{CSA}}(x):=\{y:(f(x)-y)^{\top}\Sigma^{-1}(f(x)-y)\leq\chi_{K,1-\alpha}^{2}\}$.
+Notably, if $\Sigma=\mathrm{diag}(\{\sigma_{i}^{2}\})$,
+$\mathcal{C}^{\mathrm{CSA-diag}}(x):=\left\{y:\sum^{K}_{i=1}\left(\frac{f_{k}(x)-y}{\sigma_{k}}\right)^{2}\leq\chi_{K,1-\alpha}^{2}\right\}.$ (11)
+Two insights arise from this expression.
+The first is that, unlike in the prediction regions for the case of the univariate score function $s_{k}(x,y):=(f_{k}(x)-y)^{2}$, the size $|\mathcal{C}^{\mathrm{CSA-diag}}(x)|$ is dependent on $x$.
+In the univariate case, the size is $2\widehat{q}_{k}$ across all $x$.
+Here, however, the feasible set of $y$ becomes smaller the more distinct the values $f_{k}(x)$ are.
+The second insight, therefore, is that, under such independence of residuals, prediction region sizes are minimized in having well-separated predictions, which suggests that efficiency is optimized by having an ensemble of predictors that learn distinct maps from $\mathcal{X}\to\mathcal{Y}$, such as those that focus on distinct views of the covariate space.
+Appendix F Full ImageNet Results
+Here we present the complete collection of results for the classification task across additional $\alpha$ than those that could fit in the main paper.
+Table 5: Average coverages across different coverage levels are shown in the top rows and average prediction set sizes in the bottom rows.
+Both were assessed over a batch of i.i.d.
+test samples (15% of the validation set from ImageNet).
+Standard deviations and means were computed across 10 randomized draws of the calibration and test sets.
+Dataset/$\alpha$ Metrics ResNet VGG DenseNet VFCP $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
+ImageNet Coverage 0.97 (0.011) 0.966 (0.011) 0.939 (0.002) 0.954 (0.006) 0.956 (0.011) 0.948 (0.01) 0.96 (0.013) 0.95 (0.006) 0.955 (0.013) 0.957 (0.01)
+$(\alpha=0.07)$ Length 174.828 (2.037) 181.58 (2.459) 160.623 (2.828) 61.247 (1.638) 122.899 (2.154) 111.602 (2.155) 173.264 (2.674) 86.955 (1.886) 44.787 (1.198) 45.352 (1.56)
+Coverage 0.95 (0.003) 0.949 (0.004) 0.952 (0.002) 0.95 (0.003) 0.975 (0.002) 0.954 (0.004) 0.95 (0.003) 0.949 (0.002) 0.95 (0.002) 0.95 (0.003)
+$(\alpha=0.05)$ Length 220.022 (2.072) 229.523 (3.076) 208.658 (2.016) 78.108 (2.004) 166.933 (2.157) 143.323 (2.932) 220.491 (2.773) 112.161 (2.115) 58.424 (1.674) 59.574 (3.382)
+Coverage 0.98 (0.002) 0.981 (0.002) 0.98 (0.002) 0.98 (0.002) 0.992 (0.002) 0.982 (0.002) 0.98 (0.002) 0.98 (0.003) 0.98 (0.001) 0.979 (0.002)
+$(\alpha=0.02)$ Length 357.487 (5.046) 363.376 (4.916) 342.479 (5.603) 137.356 (3.595) 311.521 (4.053) 327.754 (6.251) 355.701 (4.423) 202.573 (2.948) 117.272 (5.015) 121.477 (19.719)
+Coverage 0.99 (0.001) 0.991 (0.001) 0.989 (0.002) 0.99 (0.001) 0.997 (0.001) 0.991 (0.002) 0.99 (0.002) 0.99 (0.002) 0.99 (0.001) 0.99 (0.002)
+$(\alpha=0.01)$ Length 491.952 (6.353) 726.028 (12.157) 459.399 (6.739) 194.691 (4.579) 580.592 (7.715) 532.155 (24.829) 559.188 (7.07) 299.453 (6.526) 180.534 (8.468) 201.32 (46.509)
+Appendix G Full OpenML Results
+Table 6: Average coverages across tasks for $\alpha=0.05$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d.
+test samples (20% of the dataset size).
+Standard deviations and means were computed across 5 randomizations of draws of the training, calibration, and test sets.
+In cases where the method failed to achieve sufficient coverage (defined as $<0.93$), we do not include it in comparison for set length.
+Similarly, the single-stage approach fails to achieve coverage due to lack of exchangeability with test points.
+Dataset Metrics Linear Model LASSO Random Forest XGBoost $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
+361234 Coverage 0.97 (0.011) 0.966 (0.011) 0.939 (0.002) 0.954 (0.006) 0.956 (0.011) 0.948 (0.01) 0.96 (0.013) 0.95 (0.006) 0.955 (0.013) 0.957 (0.01)
+Length 9.673 (0.160) 9.645 (0.154) 10.080 (0.160) 9.157 (0.052) 9.196 (0.123) 8.703 (0.086) 9.524 (0.056) 17.759 (0.275) 7.646 (0.073) 7.688 (0.181)
+361235 Coverage 0.947 (0.0) 0.945 (0.005) 0.968 (0.016) 0.95 (0.005) 0.955 (0.016) 0.897 (0.005) 0.953 (0.011) 0.932 (0.021) 0.745 (0.011) 0.984 (0.005)
+Length 20.961 (0.651) 24.241 (0.246) 10.096 (0.587) 11.387 (0.452) 11.782 (0.057) 16.088 (0.118) 15.823 (1.272) 6.162 (0.458) 11.695 (0.266)
+361236 Coverage 0.975 (0.008) 0.975 (0.008) 0.961 (0.0) 0.948 (0.012) 0.948 (0.012) 0.938 (0.012) 0.965 (0.008) 0.934 (0.004) 0.94 (0.004) 0.963 (0.004)
+Length 44407.071 (1173.758) 44509.269 (1229.817) 50820.568 (385.951) 41045.069 (1221.808) 43185.942 (1002.516) 40905.295 (1089.411) 44437.938 (851.862) 60509.250 (2410.320) 30953.589 (2482.065) 33439.322 (1275.213)
+361237 Coverage 0.969 (0.023) 0.969 (0.023) 0.981 (0.0) 0.923 (0.0) 0.954 (0.015) 0.9 (0.008) 0.969 (0.023) 0.885 (0.038) 0.8 (0.015) 0.977 (0.008)
+Length 44.019 (0.990) 44.069 (1.115) 27.035 (1.014) 26.524 (1.244) 31.967 (1.118) 14.473 (0.503) 23.145 (0.199)
+361241 Coverage 0.954 (0.001) 0.956 (0.001) 0.944 (0.005) 0.957 (0.002) 0.954 (0.002) 0.923 (0.0) 0.952 (0.0) 0.949 (0.001) 0.917 (0.006) 0.951 (0.001)
+Length 19.133 (0.062) 20.245 (0.095) 18.102 (0.055) 18.482 (0.062) 17.958 (0.062) 18.932 (0.034) 29.548 (0.191) 15.199 (0.427) 17.328 (0.097)
+361242 Coverage 0.944 (0.004) 0.955 (0.0) 0.947 (0.004) 0.942 (0.0) 0.948 (0.0) 0.914 (0.003) 0.944 (0.001) 0.949 (0.003) 0.9 (0.006) 0.944 (0.002)
+Length 70.248 (0.304) 84.510 (0.282) 50.442 (0.421) 54.844 (0.036) 54.217 (0.115) 65.635 (0.094) 61.613 (0.372) 44.602 (0.170) 57.935 (0.070)
+361243 Coverage 0.922 (0.03) 0.952 (0.015) 0.956 (0.022) 0.952 (0.015) 0.937 (0.022) 0.893 (0.044) 0.937 (0.022) 0.919 (0.022) 0.748 (0.126) 0.956 (0.022)
+Length 71.388 (0.152) 75.924 (2.291) 72.877 (0.729) 68.493 (0.993) 72.048 (0.024) 43.742 (10.285) 68.220 (1.605)
+361244 Coverage 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.974 (0.015) 0.963 (0.037) 0.956 (0.015)
+Length 3.274 (0.004) 3.274 (0.004) 3.336 (0.023) 3.284 (0.010) 3.272 (0.000) 3.269 (0.003) 3.289 (0.002) 4.854 (0.293) 0.287 (0.008) 0.287 (0.008)
+361247 Coverage 0.96 (0.001) 0.953 (0.003) 0.94 (0.001) 0.951 (0.003) 0.963 (0.001) 0.903 (0.007) 0.951 (0.0) 0.954 (0.001) 0.843 (0.003) 0.943 (0.006)
+Length 0.025 (0.000) 0.038 (0.000) 0.006 (0.000) 0.016 (0.000) 0.015 (0.000) 0.022 (0.000) 0.013 (0.000) 0.005 (0.000) 0.008 (0.000)
+361249 Coverage 0.96 (0.002) 0.956 (0.002) 0.962 (0.003) 0.972 (0.007) 0.953 (0.005) 0.938 (0.002) 0.965 (0.005) 0.936 (0.01) 0.931 (0.0) 0.953 (0.005)
+Length 3.008 (0.006) 3.068 (0.009) 2.800 (0.000) 2.780 (0.025) 2.775 (0.006) 2.558 (0.019) 2.894 (0.000) 4.706 (0.099) 2.216 (0.020) 2.614 (0.043)
+Table 7: Average coverages across tasks for $\alpha=0.025$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d.
+test samples (20% of the dataset size).
+Standard deviations and means were computed across 5 randomizations of draws of the training, calibration, and test sets.
+In cases where the method failed to achieve sufficient coverage (defined as $<0.96$), we do not include it in comparison for set length.
+Similarly, the single-stage approach fails to achieve coverage due to lack of exchangeability with test points.
+Dataset Metrics Linear Model LASSO Random Forest XGBoost $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
+361234 Coverage 0.987 (0.008) 0.987 (0.008) 0.974 (0.004) 0.977 (0.008) 0.982 (0.008) 0.971 (0.01) 0.981 (0.01) 0.97 (0.008) 0.976 (0.01) 0.973 (0.006)
+Length 11.939 (0.137) 11.871 (0.084) 12.484 (0.168) 11.972 (0.009) 11.587 (0.110) 11.157 (0.086) 11.965 (0.050) 25.598 (0.974) 9.306 (0.259) 8.855 (0.059)
+361235 Coverage 0.987 (0.0) 0.982 (0.011) 0.979 (0.011) 0.984 (0.005) 0.989 (0.005) 0.966 (0.016) 0.976 (0.005) 0.958 (0.021) 0.889 (0.011) 0.989 (0.005)
+Length 24.595 (0.825) 28.841 (1.129) 11.811 (0.992) 14.237 (0.786) 14.472 (0.172) 12.278 (0.026) 19.231 (0.356) 7.719 (0.467) 12.563 (0.766)
+361236 Coverage 0.992 (0.004) 0.992 (0.004) 0.981 (0.0) 0.965 (0.008) 0.975 (0.008) 0.965 (0.008) 0.977 (0.012) 0.955 (0.008) 0.955 (0.012) 0.973 (0.004)
+Length 48591.496 (873.946) 48578.821 (867.718) 56132.760 (368.832) 46623.663 (1565.744) 47630.890 (810.373) 45714.911 (1062.420) 49188.464 (753.205) 32881.580 (3110.734) 35777.096 (1949.538)
+361237 Coverage 0.981 (0.0) 0.981 (0.0) 0.981 (0.0) 0.977 (0.008) 0.962 (0.0) 0.962 (0.0) 0.977 (0.008) 0.965 (0.008) 0.927 (0.031) 0.981 (0.0)
+Length 47.738 (0.542) 47.440 (0.959) 30.785 (0.037) 26.208 (0.897) 30.554 (0.561) 27.182 (0.803) 35.982 (0.619) 67.660 (6.380) 18.214 (0.436) 26.897 (0.515)
+361241 Coverage 0.979 (0.001) 0.978 (0.001) 0.976 (0.001) 0.978 (0.001) 0.978 (0.0) 0.964 (0.002) 0.977 (0.0) 0.972 (0.002) 0.958 (0.003) 0.979 (0.0)
+Length 21.772 (0.085) 23.089 (0.106) 21.543 (0.009) 21.454 (0.109) 20.862 (0.088) 19.291 (0.060) 21.905 (0.041) 40.082 (0.045) 17.765 (0.329) 19.897 (0.062)
+361242 Coverage 0.977 (0.003) 0.978 (0.001) 0.975 (0.002) 0.968 (0.003) 0.973 (0.004) 0.955 (0.002) 0.971 (0.002) 0.975 (0.0) 0.936 (0.001) 0.977 (0.001)
+Length 83.892 (0.143) 99.811 (0.866) 65.672 (0.212) 68.119 (0.187) 68.155 (0.357) 80.032 (0.354) 85.678 (0.371) 53.549 (0.585) 69.388 (0.128)
+361243 Coverage 0.985 (0.007) 0.985 (0.007) 0.985 (0.007) 0.956 (0.022) 0.985 (0.007) 0.97 (0.015) 0.985 (0.007) 0.978 (0.007) 0.748 (0.126) 0.985 (0.007)
+Length 92.698 (1.567) 87.949 (0.147) 88.993 (2.345) 84.569 (0.557) 79.879 (0.739) 87.950 (0.308) 137.673 (15.214) 46.504 (12.957) 79.976 (2.585)
+361244 Coverage 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.989 (0.022) 0.963 (0.037) 0.974 (0.015)
+Length 5.274 (0.004) 5.274 (0.004) 5.336 (0.023) 5.284 (0.010) 5.272 (0.000) 5.269 (0.003) 5.289 (0.002) 11.283 (1.039) 0.287 (0.008) 0.287 (0.008)
+361247 Coverage 0.98 (0.0) 0.976 (0.003) 0.969 (0.004) 0.974 (0.001) 0.977 (0.003) 0.945 (0.004) 0.971 (0.003) 0.982 (0.001) 0.906 (0.003) 0.974 (0.004)
+Length 0.029 (0.000) 0.042 (0.000) 0.009 (0.000) 0.019 (0.000) 0.018 (0.000) 0.025 (0.000) 0.016 (0.000) 0.008 (0.000) 0.012 (0.000)
+361249 Coverage 0.981 (0.003) 0.981 (0.003) 0.984 (0.002) 0.991 (0.002) 0.981 (0.003) 0.976 (0.002) 0.981 (0.003) 0.971 (0.007) 0.962 (0.011) 0.977 (0.003)
+Length 3.645 (0.023) 3.674 (0.026) 3.600 (0.000) 3.322 (0.019) 3.402 (0.005) 3.201 (0.019) 3.543 (0.000) 6.533 (0.252) 2.719 (0.164) 2.972 (0.064)
+Appendix H CSA Prediction Region Visualizations
+We now visualize some of the prediction regions corresponding to some of the trials run in Appendix G.
+While we find these intervals to be connected across these tasks, we expect visualizations over multivariate output spaces, i.e.
+for 2D regression problems, would reveal sets to be non-connected.
+Figure 9: Prediction regions across methods for task 361235 for $\alpha=0.05$ (left) and $0.025$ (right).
+Figure 10: Prediction regions across methods for task 361242 for $\alpha=0.05$ (left) and $0.025$ (right).
+Appendix I UCI Results
+We consider those regression tasks from the UCI repository [44] that have at least 1,000 samples.
 The complete collection of results is presented in Table 8.
-As in the main text, we see reduced conservatism of the CSA prediction regions compared to alternatives.
-We additionally notice that the single-stage CSA approach fails to retain the coverage guarantees.
+As discussed in the main text, across nearly all the UCI benchmark tasks, we find that the conformalized random forest does optimally and that ensembling methods provide no further benefit over simply taking this single predictor.
+In this degenerate case, we would expect an optimal aggregator to simply then return this optimal single predictor.
+We then find CSA to consistently significantly outperform other aggregation strategies and to return a prediction region of size comparable to that of the nominal random forest.
 Table 8: Average coverages across tasks for $\alpha=0.05$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d.
 test samples (10% of the dataset size).
-Standard deviations and means were computed across 10 randomizations of draws of the training, calibration, and test sets.
+We are highlighting the robustness compared to other aggregation strategies here and so bold the best performing amongst the aggregation methods.
+Standard deviations and means were computed across 5 randomizations of draws of the training, calibration, and test sets.
 Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, therefore, precluded from comparison.
-Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
-3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
-Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
-airfoil Coverage 0.95 (0.012) 0.981 (0.018) 0.954 (0.02) 0.969 (0.014) 0.962 (0.014) 0.939 (0.024) 0.969 (0.014) 0.854 (0.04) 0.952 (0.006)
-Length 18.742 (0.313) 25.218 (1.05) 7.353 (0.672) 25.256 (1.424) 18.097 (0.565) 11.96 (0.634) 18.303 (0.787) 1.873 (0.21) 2.714 (0.141)
-bike Coverage 0.963 (0.004) 0.959 (0.004) 0.948 (0.001) 0.955 (0.003) 0.955 (0.003) 0.932 (0.008) 0.959 (0.003) 0.946 (0.005) 0.955 (0.003)
-Length 3.397 (0.131) 3.416 (0.123) 0.025 (0.0) 1.716 (0.104) 1.654 (0.11) 0.837 (0.055) 2.094 (0.09) 0.569 (0.097) 0.629 (0.105)
-parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
-Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
-pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
-Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
-pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
-Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
-slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
-Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
-Appendix E Robust Traffic Routing Setup
+Dataset Metrics Linear Model LASSO Random Forest XGBoost $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
+airfoil Coverage 0.966 (0.011) 0.926 (0.011) 0.934 (0.026) 0.966 (0.011) 0.945 (0.021) 0.916 (0.016) 0.934 (0.026) 0.958 (0.032) 0.805 (0.058) 0.953 (0.011)
+Length 19.062 (0.615) 11.368 (0.188) 11.770 (0.261) 12.371 (0.131) 16.227 (0.025) 16.097 (1.052) 8.609 (0.719) 14.075 (0.734)
+bike Coverage 0.944 (0.007) 0.947 (0.004) 0.954 (0.002) 0.958 (0.003) 0.938 (0.006) 0.908 (0.0) 0.946 (0.003) 0.955 (0.0) 0.963 (0.007) 0.947 (0.003)
+Length 3.178 (0.011) 3.343 (0.021) 0.065 (0.000) 0.111 (0.000) 0.111 (0.001) 1.722 (0.007) 0.682 (0.004) 0.156 (0.019) 0.134 (0.007)
+concrete Coverage 0.977 (0.008) 0.977 (0.008) 0.981 (0.0) 0.915 (0.023) 0.962 (0.0) 0.915 (0.023) 0.981 (0.0) 0.915 (0.023) 0.854 (0.054) 0.977 (0.008)
+Length 44.295 (0.424) 44.470 (0.326) 26.053 (2.114) 26.488 (1.141) 32.455 (1.165) 19.712 (6.592) 25.302 (3.244)
+kin40k Coverage 0.949 (0.002) 0.949 (0.001) 0.946 (0.002) 0.948 (0.003) 0.945 (0.002) 0.919 (0.002) 0.949 (0.0) 0.945 (0.002) 0.907 (0.004) 0.941 (0.006)
+Length 3.781 (0.017) 3.781 (0.016) 2.333 (0.026) 3.343 (0.026) 3.269 (0.018) 3.291 (0.009) 4.985 (0.003) 2.138 (0.001) 2.456 (0.042)
+parkinsons Coverage 0.937 (0.003) 0.946 (0.0) 0.958 (0.009) 0.953 (0.007) 0.936 (0.001) 0.904 (0.008) 0.936 (0.005) 0.955 (0.004) 0.873 (0.043) 0.951 (0.003)
+Length 35.957 (0.323) 36.430 (0.328) 3.254 (0.423) 11.268 (0.114) 10.992 (0.074) 21.470 (0.003) 14.161 (0.083) 3.457 (0.727) 4.584 (0.637)
+pol Coverage 0.944 (0.001) 0.942 (0.002) 0.951 (0.004) 0.955 (0.001) 0.938 (0.001) 0.909 (0.003) 0.946 (0.001) 0.953 (0.005) 0.884 (0.009) 0.952 (0.003)
+Length 97.944 (0.150) 97.771 (0.386) 28.000 (0.000) 48.572 (1.279) 45.056 (0.821) 69.078 (0.362) 57.432 (0.663) 24.321 (1.370) 33.230 (1.569)
+protein Coverage 0.958 (0.001) 0.957 (0.002) 0.953 (0.003) 0.959 (0.003) 0.957 (0.003) 0.928 (0.003) 0.953 (0.003) 0.955 (0.0) 0.91 (0.002) 0.963 (0.004)
+Length 2.316 (0.000) 2.412 (0.011) 2.151 (0.014) 2.210 (0.006) 2.134 (0.002) 2.269 (0.001) 3.707 (0.039) 1.717 (0.036) 1.994 (0.000)
+pumadyn32nm Coverage 0.961 (0.011) 0.961 (0.01) 0.947 (0.001) 0.962 (0.003) 0.96 (0.007) 0.935 (0.003) 0.95 (0.013) 0.957 (0.013) 0.906 (0.026) 0.963 (0.001)
+Length 3.997 (0.024) 3.979 (0.031) 1.525 (0.027) 3.518 (0.058) 3.507 (0.051) 2.350 (0.043) 3.242 (0.033) 5.351 (0.139) 1.564 (0.048) 1.858 (0.078)
+tamielectric Coverage 0.953 (0.002) 0.953 (0.002) 0.947 (0.003) 0.953 (0.003) 0.952 (0.003) 0.926 (0.007) 0.949 (0.0) 0.948 (0.003) 0.909 (0.003) 0.949 (0.002)
+Length 0.950 (0.001) 0.951 (0.001) 1.271 (0.006) 0.953 (0.001) 0.948 (0.001) 1.029 (0.003) 4.539 (0.038) 0.774 (0.005) 0.799 (0.004)
+wine Coverage 0.96 (0.005) 0.943 (0.01) 0.931 (0.015) 0.923 (0.02) 0.928 (0.005) 0.884 (0.015) 0.948 (0.005) 0.946 (0.015) 0.805 (0.054) 0.933 (0.015)
+Length 2.352 (0.002) 3.521 (0.025) 2.619 (0.050) 2.621 (0.039) 3.390 (0.042) 1.683 (0.224) 2.291 (0.026)
+Appendix J Robust Traffic Routing Setup
 We replicate the experimental setup of [36], namely where a graph of Manhattan with corresponding nominal transit times was extracted using OSMnx [48].
@@
 $c_{e}:=\widetilde{c}_{e}\cdot\exp\left\{\frac{Y_{p^{e_{s}}_{x},p^{v}_{y}}+Y_{p^{e_{t}}_{x},p^{e_{t}}_{y}}}{2}\right\}.$ (12)
-Appendix F Conformal Aggregation Methods
+Appendix K Compute Details
+All OpenML were all run on a standard-grade CPU.
+The deep learning-based experiments, namely the ImageNet classification and traffic forecasting predict-then-optimize task, were performed on an Nvidia RTX 2080 Ti GPU.
+Such experiments, however, were conducted with publicly available, pre-trained models provided by the works respectively referenced in the sections describing the experimental setups.
+Appendix L Conformal Aggregation Methods
 We now describe the methods from [25] that were compared against experimentally, specifically the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches.
-As discussed in Section 2.4, these methods all follow the structural form of Equation 2 and largely differ in their choice of weights and thresholds.
-The standard majority-vote $\mathcal{C}^{M}$ is the most natural choice, defined by
+As discussed in Section 2.4, these methods all follow the structural form of
+$\mathcal{C}(x):=\left\{y\mid\sum_{k=1}^{K}w_{k}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]\geq\widehat{a}\right\}$ (13)
+and largely differ in their choice of weights and thresholds. The standard majority-vote $\mathcal{C}^{M}$ is the most natural choice, defined by
 $\mathcal{C}^{M}(x):=\left\{y\mid\frac{1}{K}\sum_{k=1}^{K}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]>\frac{1}{2}\right\}.$ (14)